### v1

# Extending Pólya’s random walker beyond probability I. Complex weights

Abstract
Working in combinatorial model $\mathrm{W_{co}}(d)$, $d=1,2,\dots$, of Pólya’s random walker in $\mathbb{Z}^{d}$, we prove two theorems on recurrence to a vertex. We obtain an effective version of the first theorem in case $d=2$. Using a semi-formal approach to generating functions, we extend both theorems beyond probability to a more general model $\mathrm{W_{\mathbb{C}}}$ with complex weights. We relate models $\mathrm{W_{co}}(d)$ to standard models $\mathrm{W_{Ma}}(d)$ based on Markov chains. The follow-up article will treat non-Archimedean models $\mathrm{W_{fo}}(k)$ in which weights are formal power series in $\mathbb{C}[[x_{1},x_{2},\dots,x_{k}]]$.
1 Introduction. Pólya’s walkers
In 1921 G. Pólya investigated in [23] particles that randomly move along the edges of the nearest neighbor graph $\mathbb{Z}^{d}$. We call this problem Pólya’s random walker. It is modeled probabilistically by Markov chains. In our article we propose and investigate alternative (combinatorial) models.
In this Section 1 we state in the language of combinatorial models $\mathrm{W_{co}}(d)$, $d\in\mathbb{N}$, Theorems 1.1 and 1.2 concerning Pólya’s random walker. They determine probabilities of recurrence to a vertex $\overline{v}$ of walks in $\mathbb{Z}^{d}$ starting at $\overline{0}$. In Theorem 1.1 and 1.2 we have $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$, respectively. We easily recognize every model $\mathrm{W_{co}}(d)$ as an element, in fact as many elements, of the more general model $\mathrm{W_{\mathbb{C}}}$. This model in Definition 1.4 goes beyond probability because it uses edge weights with values in $\mathbb{C}$. In conclusion of the section we recall standard models $\mathrm{W_{Ma}}(d)$ based on Markov chains.
In Section 2 we relate $\mathrm{W_{Ma}}(d)$ to $\mathrm{W_{co}}(d)$. Section 3 contains proofs of stronger forms of Theorems 1.1 and 1.2. In Section 4 we work out an effective version of Theorem 1.1 for $d=2$. Using our semi-formal approach to generating functions, which extends the symbolic method in combinatorics [3, 8, 10, 12, 28] from finite to countable sets, in Section 5 we extend (strengthened) Theorems 1.1 and 1.2 to $\mathrm{W_{\mathbb{C}}}$. In Section 6 we outline the follow-up article which will deal with non-Archimedean weights. At the very end of this section, we preview our results in more detail.
G. Pólya extended his pioneering article [23] in [24]; see [4] for his scientific biography. Pólya’s random walker in $\mathbb{Z}^{d}$ is discussed, for example, in [5, 6, 7, 9, 11, 13, 17, 19, 21, 22, 25, 26, 30, 31].
We review notation. =$\mathbb{N}=\{1,2,\dots\}$, $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ and $\mathbb{Z}$ are the integers. $\mathbb{R}$ and $\mathbb{C}$ denote the fields of real and complex numbers, and $\mathbb{R}_{\geq 0}$ are nonnegative real numbers. For a set $A$ we denote by $\binom{A}{2}$ the set of two-element subsets of $A$. Instead of $\binom{\mathbb{N}}{2}$ we write $\mathbb{N}_{2}$. If $A$ is a finite set, $|A|$ ($\in\mathbb{N}_{0}$) denotes the number of its elements. $\mathcal{P}(A)=\{X\colon\;X\subset A\}$ is the power set of $A$. For $n\in\mathbb{N}$ let $[n]=\{1,2,\dots,n\}$, and let $[0]=\emptyset$. If $f\colon X\to Y$ is a map and $Z$ is any set, then $f[Z]=\{f(x)\colon\;x\in X\cap Z\}$ ($\subset Y$) and $f^{-1}[Z]=\{x\in X\colon\;f(x)\in Z\}$ ($\subset X$).
Let $d\in\mathbb{N}$. The nearest neighbor graph on $\mathbb{Z}^{d}$ is the countable graph
${\textstyle G_{d}=(\mathbb{Z}^{d},\,E_{d}),\ E_{d}\subset\binom{\mathbb{Z}^{d}}{2}\,,}$
with the vertices $\mathbb{Z}^{d}=\mathbb{Z}\times\mathbb{Z}\times\dots\times\mathbb{Z}$ ($d$ factors) and edges
${\textstyle\{\overline{u},\,\overline{v}\}=\{\langle u_{1},\,u_{2},\,\dots,\,u_{d}\rangle,\,\langle v_{1},\,v_{2},\,\dots,\,v_{d}\rangle\}\in E_{d}\iff\sum_{i=1}^{d}|u_{i}-v_{i}|=1\,.}$
$G_{d}$ has no multiple edges and no loops, and is $2d$-regular (every vertex lies in $2d$ edges). It is vertex-transitive.
We consider walks in $G_{d}$ starting at the vertex 0¯=($\overline{0}=(0,0\dots,0)$. A walk in $G_{d}$ is any tuple
$w=\langle\overline{u}_{0},\,\overline{u}_{1},\,\dots,\,\overline{u}_{n}\rangle,\ n\in\mathbb{N}_{0}\,,$
of vertices $\overline{u}_{i}\in\mathbb{Z}^{d}$ such that $\{\overline{u}_{i-1},\overline{u}_{i}\}\in E_{d}$ for every $i\in[n]$. The set $\{\overline{u}_{1},\overline{u}_{2},\dots,\overline{u}_{n-1}\}$ is the inner of $w$. The length $n$ of $w$ is denoted by $|w|$. We say that $w$ starts at $\overline{u}_{0}$, or that it is a $\overline{u}_{0}$-walk, and that it ends at $\overline{u}_{n}$. We call $w$ recurrent if it starts at $\overline{0}$, $|w|=n>0$ and $\overline{u}_{i}=\overline{0}$ for some $i\in[n]$. Thus $w$ revisits its start $\overline{0}$. Due to the vertex-transitivity of $G_{d}$, the starting vertex $\overline{0}$ may be replaced with any other vertex. For $d\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$ we let $W_{d}(n)$ denote the set of $\overline{0}$-walks of length $n$ in $G_{d}$. Thus $|W_{d}(n)|=(2d)^{n}$. We denote by $W_{d,\mathrm{rec}}(n)$ ($\subset W_{d}(n)$) the set of recurrent walks of length $n$. Theorem 1.1 below is a combinatorial version of the classical result on random walks in $G_{d}$. It is not Pólya’s theorem which comes in Theorem 1.2.
Theorem 1.1 ($\overline{0}$-recurrence)
For every $d\in\mathbb{N}$,
$\lim_{n\to\infty}\frac{|W_{d,\mathrm{rec}}(\,n)|}{|W_{d}(n)|}=\lim_{n\to\infty}\frac{|W_{d,\mathrm{rec}}(\,n)|}{(2d)^{n}}\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\;.\end{array}\right.$
So for $d\leq 2$ and $n\to\infty$ the proportion of non-recurrent walks goes to 00. For $d\geq 3$ and every $n\in\mathbb{N}_{0}$ it is greater than $c>0$. The strengthening in Theorem 3.9 gives a formula for the limit if $d\geq 3$.
For $d\in\mathbb{N}$, $\overline{v}\in\mathbb{Z}^{d}$ and $n\in\mathbb{N}_{0}$ let $W_{d}(\overline{v},n)$ ($\subset W_{d}(n)$) be the set of $\overline{0}$-walks of length $n$ in $G_{d}$ such that $\overline{u}_{i}=\overline{v}$ for some $i\in[n]$. So $W_{d,\mathrm{rec}}(n)=W_{d}(\overline{0},n)$. We call walks in $W_{d}(\overline{v},n)$ $\overline{v}$-recurrent. The theorem below is a combinatorial version of Pólya’s classical result in [23].
Theorem 1.2 ($\overline{v}$-recurrence)
For every $d\in\mathbb{N}$ and every $\overline{v}\in\mathbb{Z}^{d}$,
$\lim_{n\to\infty}\frac{|W_{d}(\overline{v},\,n)|}{|W_{d}(n)|}=\lim_{n\to\infty}\frac{|W_{d}(\overline{v},\,n)|}{(2d)^{n}}\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\;.\end{array}\right.$
By [20, Theorem 1.5.4] and our Corollary 2.5, for every $d\geq 3$ the limit is asymptotic to
$\frac{\frac{d}{2}\Gamma(\frac{d}{2}-1)\pi^{-d/2}}{\|\overline{v}\|^{d-2}}\ \ \ \ (\|\overline{v}\|\to+\infty)\,,$
where $\Gamma(x)=\int_{0}^{+\infty}t^{x-1}\mathrm{e}^{-t}\,\mathrm{d}t$ is the Gamma function and
${\textstyle\|\overline{v}\|=\sqrt{v_{1}^{2}+v_{2}^{2}+\dots+v_{d}^{2}}}$
is the Euclidean norm. The strengthening in Theorem 3.11 gives a formula for the limit if $d\geq 3$.
In [23], see [23, top of p. 157], Pólya proved Theorem 1.2. He announces it in [23, pp. 149–150].
For concreteness we imagine that the wandering point starts its journey at the origin of coordinates at the time $t=0$ and that it moves with speed $1$.
[]delimited-[][$\dots$][ … ]
Now I treat the wandering in the $d$-dimensional net of lines; a node with the coordinates $a_{1},a_{2},\dots,a_{d}$ is given; in question is the probability that in the time span $0<t\leq n$ the wandering point coincides at least once with the given node $a_{1},a_{2},\dots,a_{d}$. The probability obviously grows with $n$. A question arises: does it tend to certainty when $n$ grows without limit?
Yes when $d=1$ or $d=2$, no when $d\geq 3$. In the following, I will justify this answer. [our translation, original emphasize]
Pólya’s results are interpreted in terms of (modern) Markov chains, but this is an anachronism. Although A. Markov first published on them in 1906, rigorous probability theory was born in [18] one decade after [23].
In order to introduce the model $\mathrm{W_{\mathbb{C}}}$ we review complex series. Let $X$ be an at most countable set and $g\colon X\to\mathbb{C}$ be a map. We call $g$series (on $X$) and denote it by $\sum_{x\in X}g(x)$. We say that $\sum_{x\in X}g(x)$ absolutely converges if two equivalent conditions hold.
  1. 1.
    There is a constant $c>0$ such that for every finite set $Y\subset X$ we have $\sum_{x\in Y}|g(x)|\leq c$.
  2. 2.
    If $X$ is countable then for every bijection $f\colon\mathbb{N}\to X$ the limit
    ${\textstyle s=\lim_{n\to\infty}\sum_{i=1}^{n}g(f(i))\ \ (\in\mathbb{C})}$
    exists and is independent of $f$.
We call $s$ the sum of the series and denote it also by $\sum_{x\in X}g(x)$. For a finite set $X=\{x_{1},x_{2},\dots,x_{n}\}$ the series always absolutely converges and the sum is just
$g(x_{1})+g(x_{2})+\dots+g(x_{n})\,.$
If $S=\sum_{x\in X}g(x)$ is a series and $Y\subset X$, we call $\sum_{x\in Y}g(x)$subseries of $S$. It is easy to see that any subseries of an absolutely convergent series absolutely converges. If $X$ is countable, $g\colon X\to\mathbb{R}$ and for every bijection $f\colon\mathbb{N}\to X$ we have $\lim_{n\to\infty}\sum_{i=1}^{n}g(f(i))=+\infty$, we say that the series $\sum_{x\in X}g(x)$ has sum +$\infty$+ ∞. If $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$ is a formal power series, or FPS, in $\mathbb{C}[[x]]$, we denote the function given by sums of $U(x)$ for $x\in\mathbb{C}$ by $F_{U}(x)$.
Let $n\in\mathbb{N}_{0}$ and $K_{\mathbb{N}}=(\mathbb{N},\binom{\mathbb{N}}{2})=(\mathbb{N},\mathbb{N}_{2})$ be the complete countable graph. A walk with length $n$ in $K_{\mathbb{N}}$ is any tuple
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{n}\rangle,\ n\in\mathbb{N}_{0}\,,$
of vertices $u_{i}\in\mathbb{N}$ such that $u_{i-1}\neq u_{i}$ for every $i\in[n]$. We denote the length $n$ of $w$ by $|w|$. We say that $w$ starts at $u_{0}$, or that it is a $u_{0}$-walk, and that it ends at $u_{n}$. The set $\{u_{1},u_{2},\dots,u_{n-1}\}$ is the inner of $w$. We consider walks starting at $1$. For $n\in\mathbb{N}_{0}$, let
$W(n)=\{w\colon\;\text{$w$ is a~{}$1$-walk in $K_{\mathbb{N}}$ with $|w|=n$}\}\;.$
Definition 1.3 ($G(h)$)
Let $X$ be a set with a special element $0_{X}\in X$ and let $h\colon\mathbb{N}_{2}\to X$. We define the graph
$G(h)=(V(h),\,E(h))$
as the connected component containing $1$ of the subgraph of $K_{\mathbb{N}}$ formed by the edges $e$ with $h(e)\neq 0_{X}$.
If $h\colon\mathbb{N}_{2}\to X$, where
[…]
is a commutative ring, we extend $h$ to any walk $w=\langle u_{0},u_{1},\dots,u_{n}\rangle$ in $K_{\mathbb{N}}$ as follows. For $n=0$ we set $h(w)=1_{X}$. For $n>0$ we set
${\textstyle h(w)=\prod_{i=1}^{n}h(\{u_{i-1},\,u_{i}\})\,.}$
If all vertex degrees in the graph $G(h)$ are finite, then for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ has only finitely many nonzero summands and has a sum. More generally, we call a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ light if for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges.
Definition 1.4 ($\mathrm{W_{\mathbb{C}}}$)
The model $\mathrm{W_{\mathrm{\mathbb{C}}}}$ is the set of light weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We call it Pólya’s complex walker.
In Section 5 we extend Theorems 1.1 and 1.2, in fact Theorems 3.9 and 3.11, to $\mathrm{W_{\mathrm{\mathbb{C}}}}$
We easily recognize models $\mathrm{W_{\mathrm{co}}}(d)$ as elements in $\mathrm{W_{\mathbb{C}}}$. We omit the proof of the next result.
Proposition 1.5
Let $d,n,u\in\mathbb{N}$. Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is such that $G(h)$ is isomorphic to $G_{d}$, that always $h(e)\in\{0,\frac{1}{2d}\}$, that $1$ corresponds in the isomorphism to $\overline{0}$, and $u$ to $\overline{v}$. Then we have for walks $w=\langle u_{0},u_{1},\dots,u_{n}\rangle$ in $K_{\mathbb{N}}$ that
${\textstyle(2d)^{-n}\cdot|W_{d}(\overline{v},\,n)|=\sum_{\begin{subarray}{c}w\in W(n)\\ \exists i\in[n]:\,u_{i}=u\end{subarray}}h(w)\,.}$
The proportion or, if you want, probability of $\overline{v}$-recurrent walks of length $n$ in $G_{d}$ equals to the total $h$-weight of $u$-recurrent walks of length $n$ in $K_{\mathbb{N}}$.
We review the Markov chain model $\mathrm{W_{Ma}}(d)$ of Pólya’s random walker. Recall that a probability space is a triple
$P=\langle\Omega,\,\mathcal{A},\,\mathrm{Pr}\rangle$
of a nonempty set $\Omega$, a $\sigma$-algebra $\mathcal{A}$ ($\subset\mathcal{P}(\Omega)$) on the set $\Omega$ and a $\sigma$-additive probability map $\mathrm{Pr}\colon\mathcal{A}\to[0,1]$ such that $\mathrm{Pr}(\Omega)=1$. In more details, $\mathcal{A}\neq\emptyset$, $\mathcal{A}$ is closed to complements to $\Omega$ and to countable unions, and
${\textstyle\mathrm{Pr}\big{(}\bigcup_{n=1}^{\infty}A_{n}\big{)}=\sum_{n=1}^{\infty}\mathrm{Pr}(A_{n})}$
for any mutually disjoint sets $A_{n}\in\mathcal{A}$, $n\in\mathbb{N}$. Let $A,B\in\mathcal{A}$ with $\mathrm{Pr}(B)>0$. The conditional probability of $A$ given $B$ is the ratio
${\textstyle\mathrm{Pr}(A\,|\,B)=\frac{\mathrm{Pr}(A\cap B)}{\mathrm{Pr}(B)}\,.}$
For $\mathrm{Pr}(B)=0$ this quantity is not defined. Let $S$ be an at most countable set. An $S$-valued $P$-random variable is any map $X\colon\Omega\to S$ such that $X^{-1}[\{s\}]\in\mathcal{A}$ for every $s\in S$. We denote these elements of $\mathcal{A}$ by “$X=s$”. Their intersections and unions are indicated by $\wedge$ and $\vee$, respectively.
Definition 1.6 ($\mathrm{W_{Ma}}(d)$)
Let $P=\langle\Omega,\mathcal{A},\mathrm{Pr}\rangle$ be a probability space and
$(X_{0},\,X_{1},\,X_{2},\,\,\dots)$
be an infinite sequence of $\mathbb{Z}^{d}$-valued $P$-random variables. If the following three conditions are satisfied, we call this sequence Pólya’s Markov walker and denote it $\mathrm{W_{Ma}}(d)$.
  1. 1.
    For every $n\in\mathbb{N}_{0}$ and every $\overline{u}_{0}$, $\overline{u}_{1}$, $\dots$, $\overline{u}_{n+1}$ in $\mathbb{Z}^{d}$ with $\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{n}=\overline{u}_{n})>0$ we have
    $\mathrm{Pr}(X_{n+1}=\overline{u}_{n+1}\,|\,X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{n}=\overline{u}_{n})$
    $=\mathrm{Pr}(X_{n+1}=\overline{u}_{n+1}\,|\,X_{n}=\overline{u}_{n})\,.$
  2. 2.
    $\mathrm{Pr}(X_{0}=\overline{0})=1$.
  3. 3.
    If $n\in\mathbb{N}$, $\overline{u},\overline{v}\in\mathbb{Z}^{d}$ and $\mathrm{Pr}(X_{n-1}=\overline{u})>0$, then
    $\mathrm{Pr}(X_{n}=\overline{v}\,|\,X_{n-1}=\overline{u})=\left\{\begin{array}[]{lll}{\textstyle\frac{1}{2d}}&\dots&\{\overline{u},\,\overline{v}\}\in E_{d}\text{ and}\\ 0&\dots&\{\overline{u},\,\overline{v}\}\not\in E_{d}\,.\end{array}\right.$
A preview of this article
In Section 2 we obtain six results on relations between models $\mathrm{W_{Ma}}(d)$ and $\mathrm{W_{co}}(d)$. For example, Corollary 2.5 says that for every $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$,
${\textstyle\mathrm{Pr}\big{(}\bigvee_{n=1}^{\infty}X_{n}=\overline{v}\big{)}=\lim_{n\to\infty}(2d)^{-n}\cdot|W_{d}(\overline{v},\,n)|\,.}$
In Section 3 we strengthen Theorems 1.1 and 1.2 to Theorems 3.9 and 3.11, respectively. The strengthened theorems give explicit formulas for the limits when $d\geq 3$. They are
${\textstyle 1-\frac{1}{B(1)}\,\text{ and }\,\sqrt{1-\frac{B_{0}(1)}{B(1)}}\,,}$
respectively, with $B(x)$ and $B_{0}(x)$ being generating functions of closed walks. In Section 4 we obtain an effective form of Theorem 1.1 for $d=2$. For example, Corollary 4.6 says that for every $N\geq N_{0}$, with $N_{0}$ effectively computable, we have
$(0.9\log N)^{-1}\leq 1-4^{-N}|W_{2,\mathrm{rec}}(N)|\leq(0.1\log N)^{-1}\,.$
Sections 2 and 5 contain our main results. In Section 5 on $\mathrm{W_{\mathbb{C}}}$ we generalize Theorem 3.9 to Theorems 5.18 and 5.19, and Theorem 3.11 to Theorems 5.22 and 5.23. The first theorem in each pair allows any (light and, respectively, $v$-transitive) weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and the computed main quantities are
${\textstyle\lim_{n\to\infty}\sum_{j=0}^{n}a_{j}^{h}\,\text{ and }\,\lim_{n\to\infty}\sum_{j=0}^{n}(a_{j}^{h})^{\prime}\,,}$
where $a_{j}^{h}$ and $(a_{j}^{h})^{\prime}$ are, respectively, the total $h$-weights of recurrent and $v$-recurrent walks of length $j$ in $K_{\mathbb{N}}$. The second theorem in each pair allows only convex (light and, respectively, $v$-transitive) weight $h$, but the calculated main quantities are more precise:
${\textstyle\lim_{n\to\infty}a_{n}^{h}\,\text{ and }\,\lim_{n\to\infty}(a_{n}^{h})^{\prime}\,,}$
respectively. Convexity of $h$ roughly means that always $\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})=1$.
2 Relations between models $\mathrm{W_{Ma}}(d)$ and $\mathrm{W_{co}}(d)$
We establish several relations between models $\mathrm{W_{Ma}}(d)$ and $\mathrm{W_{co}}(d)$. The next main theorem shows that Pólya’s Markov walker correctly recognizes walks among tuples of vertices in the graph $G_{d}$.
Theorem 2.1
Let $\mathrm{Pr}$ and $X_{n}$ be as in Definition 1.6. If $d\in\mathbb{N}$, $n\in\mathbb{N}_{0}$ and $t=\langle\overline{u}_{0},\overline{u}_{1},\dots,\overline{u}_{n}\rangle\in(\mathbb{Z}^{d})^{n+1}$, then
$\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{n}=\overline{u}_{n})=\left\{\begin{array}[]{lll}{\textstyle\big{(}\frac{1}{2d}\big{)}^{n}}&\dots&t\in W_{d}(n)\text{ and}\\ 0&\dots&t\not\in W_{d}(n)\,.\end{array}\right.$
Proof. Let $n\in\mathbb{N}_{0}$ and $w=\langle\overline{u}_{0},\overline{u}_{1},\dots,\overline{u}_{n}\rangle$ be any $\overline{0}$-walk in $G_{d}$. We prove by induction on $i=0,1,\dots,n$ that
$\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{i}=\overline{u}_{i})=(2d)^{-i}\,.$
For $i=0$ this holds by part 2 of Definition 1.6. Let $0<i<n$. Then the probability $\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{i+1}=\overline{u}_{i+1})$ indeed equals
$\mathrm{Pr}(X_{i+1}=\overline{u}_{i+1}\,|\,X_{0}=\overline{u}_{0}\wedge\dots\wedge X_{i}=\overline{u}_{i})\cdot\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge\dots\wedge X_{i}=\overline{u}_{i})$
$=\mathrm{Pr}(X_{i+1}=\overline{u}_{i+1}\,|\,X_{i}=\overline{u}_{i})\cdot(2d)^{-i}$
$=(2d)^{-1}\cdot(2d)^{-i}=(2d)^{-i-1}\;.$
In the first line we use the formula for the conditional probability and the inductive assumption that
$\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{i}=\overline{u}_{i})=(2d)^{-i}>0\,.$
We use this assumption also on the second line, together with part 1 of Definition 1.6. On the third line we use part 3 of Definition 1.6 and the fact that $\{\overline{u}_{i},\overline{u}_{i+1}\}\in E_{d}$.
We prove the second case. Let $n\in\mathbb{N}_{0}$ and $\langle\overline{u}_{0},\overline{u}_{1},\dots,\overline{u}_{n}\rangle$ be in $(\mathbb{Z}^{d})^{n+1}\setminus W_{d}(n)$. Let $j\in[n-1]\cup\{0\}$ be maximum such that
$\langle\overline{u}_{0},\,\overline{u}_{1},\,\dots,\,\overline{u}_{j}\rangle\in W_{d}(j)\;.$
If $j$ does not exist, that is if $\overline{u}_{0}\neq\overline{0}$, we set $j=-1$. By the monotonicity of probability,
$\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge\dots\wedge X_{n}=\overline{u}_{n})\leq\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge\dots\wedge X_{j+1}=\overline{u}_{j+1})\;.$
If $j=-1$, the last probability is $\mathrm{Pr}(X_{0}=\overline{u}_{0})=0$ by part 2 of Definition 1.6 because $\overline{u}_{0}\neq\overline{0}$. For $j\geq 0$ we set $C=(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{j}=\overline{u}_{j})$ ($\in\mathcal{A}$) and get that
$\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge\dots\wedge X_{j+1}=\overline{u}_{j+1})$ $=$ $\mathrm{Pr}(X_{j+1}=\overline{u}_{j+1}\,|\,C)\cdot\mathrm{Pr}(C)$
$=$ $\mathrm{Pr}(X_{j+1}=\overline{u}_{j+1}\,|\,X_{j}=\overline{u}_{j})\cdot(2d)^{-j}$
$=$ $0\displaystyle 0\cdot(2d)^{-j}=$ 0
— thus $\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{n}=\overline{u}_{n})=0$. In the first line we use the formula for the conditional probability and the fact proved in the first case that $\mathrm{Pr}(C)=(2d)^{-j}>0$. In the second line we use this fact too, and part 1 of Definition 1.6. In the third line we use part 3 of Definition 1.6 and the fact that $\{\overline{u}_{j},\overline{u}_{j+1}\}\not\in E_{d}$ by the maximality of $j$. $\Box$
Corollary 2.2
If $d\in\mathbb{N}$, $n\in\mathbb{N}_{0}$ and $\overline{v}\in\mathbb{Z}^{d}$, then
$\mathrm{Pr}(X_{n}=\overline{v})=(2d)^{-n}\cdot|\{w\in W_{d}(n)\colon\;\text{$w$ ends at $\overline{v}$}\}|\;.$
Proof. Let $d$, $n$ and $\overline{v}$ be as stated. If $n=0$, the equality holds by part 2 of Definition 1.6: if $\overline{v}=\overline{0}$ (respectively $\overline{v}\neq\overline{0}$) then both sides are $1$ (respectively 00). If $n\in\mathbb{N}$, we let
$\widehat{u}=\langle\overline{u}_{0},\,\overline{u}_{1},\,\dots,\,\overline{u}_{n-1}\rangle$
run in $(\mathbb{Z}^{d})^{n}$ and set $C=(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{n-1}=\overline{u}_{n-1})$ ($\in\mathcal{A}$). The $\sigma$-additivity of $\mathrm{Pr}(\cdot)$ and Theorem 2.1 give
${\textstyle\mathrm{Pr}(X_{n}=\overline{v})=\sum_{\widehat{u}}\mathrm{Pr}(C\wedge X_{n}=\overline{v})=\sum_{\widehat{u},\;\widehat{u}\,\overline{v}\in W_{d}(n)}(2d)^{-n}\,,}$
which equals to the stated product. $\Box$
For $\overline{u}=\langle u_{1},u_{2},\dots,u_{d}\rangle$ in $\mathbb{Z}^{d}$ we set $\|\overline{u}\|=|u_{1}|+|u_{2}|+\dots+|u_{d}|$.
Corollary 2.3
Let $d\in\mathbb{N}$, $m,n\in\mathbb{N}_{0}$ and $\overline{u}\in\mathbb{Z}^{d}$ with $\|\overline{u}\|=n$. Then $\mathrm{Pr}(X_{n}=\overline{u})>0$ and $\mathrm{Pr}(X_{m}=\overline{u})=0$ for $m<n$.
Proof. Let $\overline{u}\in\mathbb{Z}^{d}$ with $\|\overline{u}\|=n$ ($\in\mathbb{N}_{0}$). Corollary 2.2 gives $\mathrm{Pr}(X_{n}=\overline{u})>0$ because there exists a $\overline{0}$-walk $w=\langle\overline{u}_{0},\overline{u}_{1},\dots,\overline{u}_{n}\rangle$ in $G_{d}$ with $\overline{u}_{n}=\overline{u}$. On the other hand, if $m\in\mathbb{N}_{0}$ and $m<n=\|\overline{u}\|$, then there is no $\overline{0}$-walk of length $m$ in $G_{d}$ ending at $\overline{u}$, and Corollary 2.2 gives $\mathrm{Pr}(X_{m}=\overline{u})=0$. $\Box$
We leave it as an exercise for the reader to show that for $m>n=\|\overline{u}\|$ one has $\mathrm{Pr}(X_{m}=\overline{u})>0$ iff $m\equiv n$ modulo $2$.
Recall that $W_{d}(\overline{v},n)$ is the set of $\overline{0}$-walks of length $n$ in the graph $G_{d}$ such that they visit the vertex $\overline{v}$ again at a step $i>0$.
Corollary 2.4
If $d\in\mathbb{N}$, $n\in\mathbb{N}_{0}$ and $\overline{v}\in\mathbb{Z}^{d}$, then
${\textstyle\mathrm{Pr}\big{(}\bigvee_{i=1}^{n}X_{i}=\overline{v}\big{)}=(2d)^{-n}\cdot|W_{d}(\overline{v},\,n)|\,.}$
Proof. Let $d$, $n$ and $\overline{v}$ be as stated. For $n\in\mathbb{N}_{0}$ we let $\widehat{u}=\langle\overline{u}_{0},\overline{u}_{1},\dots,\overline{u}_{n}\rangle$ run in $(\mathbb{Z}^{d})^{n}$. By the $\sigma$-additivity of $\mathrm{Pr}(\cdot)$ and Theorem 2.1, the stated probability equals to
${\textstyle\sum_{\begin{subarray}{c}\widehat{u}\in(\mathbb{Z}^{d})^{n}\\ \exists\,i\in[n]:\,\overline{u}_{i}=\overline{v}\end{subarray}}\mathrm{Pr}(X_{0}=\overline{u}_{0}\wedge X_{1}=\overline{u}_{1}\wedge\dots\wedge X_{n}=\overline{u}_{n})=\sum_{\begin{subarray}{c}\widehat{u}\in W_{d}(n)\\ \exists\,i\in[n]:\,\overline{u}_{i}=\overline{v}\end{subarray}}(2d)^{-n}\,,}$
which equals to the stated product. $\Box$
Corollary 2.5
If $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$, then
${\textstyle\mathrm{Pr}\big{(}\bigvee_{n=1}^{\infty}X_{n}=\overline{v}\big{)}=\lim_{n\to\infty}(2d)^{-n}\cdot|W_{d}(\overline{v},\,n)|\,.}$
Proof. Let $\mathcal{A}$ be as in Definition 1.6. For every sequence of sets $A_{n}\in\mathcal{A}$, $n\in\mathbb{N}$, we have
${\textstyle\mathrm{Pr}\big{(}\bigcup_{i=1}^{\infty}A_{i}\big{)}=\lim_{n\to\infty}\mathrm{Pr}\big{(}\bigcup_{i=1}^{n}A_{i}\big{)}\,.}$
This follows from the identity $\bigcup_{i=1}^{\infty}A_{i}=\bigcup_{n=1}^{\infty}(\bigcup_{i=0}^{n}A_{i}\setminus\bigcup_{i=0}^{n-1}A_{i})$, where $A_{0}=\emptyset$, from the equality $\mathrm{Pr}(A\setminus B)=\mathrm{Pr}(A)-\mathrm{Pr}(B)$ for sets $B\subset A$ in $\mathcal{A}$, and from the $\sigma$-additivity of $\mathrm{Pr}(\cdot)$. Since by Corollary 2.4 we have for every $n\in\mathbb{N}$ that
${\textstyle\mathrm{Pr}\big{(}\bigvee_{i=1}^{n}X_{i}=\overline{v}\big{)}=(2d)^{-n}\cdot|W_{d}(\overline{v},\,n)|\,,}$
it follows that $\mathrm{Pr}\big{(}\bigvee_{n=1}^{\infty}X_{n}=\overline{v}\big{)}$ equals
${\textstyle\lim_{n\to\infty}\mathrm{Pr}\big{(}\bigvee_{i=1}^{n}X_{i}=\overline{v}\big{)}=\lim_{n\to\infty}(2d)^{-n}\cdot|W_{d}(\overline{v},\,n)|\,.}$
$\Box$
We conclude this section with a trivial situation which is of some interest. For any walk $w=\langle\overline{u}_{0},\overline{u}_{1},\dots,\overline{u}_{n}\rangle$ in $G_{d}$ we denote its last vertex $\overline{u}_{n}$ by $\ell(w)$.
Proposition 2.6
If $d\in\mathbb{N}$, $n\in\mathbb{N}_{0}$ and $\overline{u},\overline{v}\in\mathbb{Z}^{d}$ with $\overline{u}\neq\overline{v}$, then
$\mathrm{Pr}(X_{n}=\overline{u}\wedge X_{n}=\overline{v})=0=(2d)^{-n}\cdot|\{w\in W_{d}(n)\colon\;\ell(w)=\overline{u}\wedge\ell(w)=\overline{v}\}|\,.$
Proof. The latter equality to 00 is obvious. But so is the former, for $\overline{u}\neq\overline{v}$ we have
$X_{n}^{-1}[\{\overline{u}\}]\cap X_{n}^{-1}[\{\overline{v}\}]=\emptyset\,.$
$\Box$
Today the only model of Pólya’s random walker in $\mathbb{Z}^{d}$ is $\mathrm{W_{Ma}}(d)$. If $\mathrm{W_{co}}(d)$ is mentioned, it is only as an illustration or motivation. In [14] we hope to generalize the above relations between $\mathrm{W_{Ma}}(d)$ and $\mathrm{W_{co}}(d)$. Definition 1.6 is non-vacuous only if the Markov chain $\mathrm{W_{Ma}}(d)$ and the probability space $P$ supporting it exist. They indeed exist but their rigorous construction requires some effort, see [7, 32]. In contrast, the existence of sets of walks $W_{d}(n)$ and $W_{d}(\overline{v},n)$ in the model $\mathrm{W_{co}}(d)$ is obvious, without complicated justification.
3 Theorems 1.1 and 1.2 in stronger forms
We strengthen Theorem 1.1 on $\overline{0}$-recurrence to Theorem 3.9, and Theorem 1.2 on $\overline{v}$-recurrence to Theorem 3.11. In the strengthening we obtain explicit formulas for the limits if $d\geq 3$ (for $d\leq 2$ the limit is $1$).
First we strengthen Theorem 1.1. We define four sequences $(a_{n})$, $(b_{n})$, $(c_{n})$ and $(d_{n})$ in $\mathbb{N}_{0}$ counting certain $\overline{0}$-walks of length $n$ in $G_{d}$. Let $n\in\mathbb{N}_{0}$. We define $a_{n}$ as the number of recurrent walks, $b_{n}$ as the number of walks ending at $\overline{0}$, $c_{n}$ as the number of walks ending at $\overline{0}$ but with the inner avoiding $\overline{0}$, and $d_{n}=(2d)^{n}$ as the number of all walks. We have $a_{0}=c_{0}=0$ and $b_{0}=d_{0}=1$. Let the FPS
${\textstyle A(x)=\sum_{n=0}^{\infty}\frac{a_{n}}{d_{n}}x^{n},\ B(x)=\sum_{n=0}^{\infty}\frac{b_{n}}{d_{n}}x^{n},\ C(x)=\sum_{n=0}^{\infty}\frac{c_{n}}{d_{n}}x^{n}}$
and ${\textstyle\text{and }\,D(x)=\sum_{n=0}^{\infty}\frac{d_{n}}{d_{n}}x^{n}=\frac{1}{1-x}}$
in $\mathbb{C}[[x]]$ be the corresponding generating functions. The numbers $d_{n}$ are multiplicative: for every $m,n\in\mathbb{N}_{0}$ we have $d_{m+n}=d_{m}d_{n}$. We suppress in the notation the dependence on $d$. The quantities of the primary interest are $a_{n}$ and $A(x)$.
We establish for the above generating functions two relations.
Proposition 3.1
Let $A(x)$, $B(x)$, $C(x)$ and $D(x)$ be as above. They satisfy the following formal relations.
  1. 1.
    $A(x)=C(x)D(x)=C(x)\frac{1}{1-x}$.
  2. 2.
    $B(x)=\frac{1}{1-C(x)}=\sum_{j=0}^{\infty}C(x)^{j}$, equivalently $C(x)=1-\frac{1}{B(x)}$.
Proof. 1. If $n\in\mathbb{N}_{0}$ then
${\textstyle a_{n}=\sum_{j=0}^{n}c_{j}d_{n-j},\,\text{ so that }\,\frac{a_{n}}{d_{n}}=\sum_{j=0}^{n}\frac{c_{j}}{d_{j}}\cdot 1\,.}$
The first equality follows from the unique splitting of any recurrent walk $w$ of length $n$ in $G_{d}$ at the first revisit of $\overline{0}$ in step $j$ as $w=w_{1}w_{2}$. The walks $w_{1}$ are counted by $c_{j}$, and the walks $w_{2}$ by $d_{n-j}$. Dividing by $d_{n}$ and using that $d_{n}=d_{j}d_{n-j}$ we get the second equality. In $\mathbb{C}[[x]]$ it reads as $A(x)=C(x)D(x)$.
2. Now $b_{0}=1$ and for $n\in\mathbb{N}$,
${\textstyle b_{n}=\sum_{j=1}^{\infty}\sum_{\begin{subarray}{c}l_{1},\dots,l_{j}\in\mathbb{N}\\ l_{1}+\dots+l_{j}=n\end{subarray}}c_{l_{1}}c_{l_{2}}\dots c_{l_{j}}\,.}$
This equality follows from the unique decomposition of any $\overline{0}$-walk of length $n\in\mathbb{N}$ ending at $\overline{0}$ into $j\in\mathbb{N}$ walks with lengths $l_{i}$ such that each starts and ends at $\overline{0}$ and has inner avoiding $\overline{0}$. Dividing by $d_{n}$ and using the multiplicativity of $d_{n}$ we get that for every $n\in\mathbb{N}_{0}$,
${\textstyle\frac{b_{n}}{d_{n}}=\sum_{j=0}^{\infty}\sum_{\begin{subarray}{c}l_{1},\dots,l_{j}\in\mathbb{N}\\ l_{1}+\dots+l_{j}=n\end{subarray}}\frac{c_{l_{1}}}{d_{l_{1}}}\cdot\frac{c_{l_{2}}}{d_{l_{2}}}\cdot\ldots\cdot\frac{c_{l_{j}}}{d_{l_{j}}}\;.}$
Multiplying by $x^{n}$ and formally summing the results we get the relation $B(x)=\sum_{j=0}^{\infty}C(x)^{j}=\frac{1}{1-C(x)}$. Solving for $C(x)$ yields the second equality. $\Box$
We denote by $\mathbb{R}_{\geq 0}[[x]]$ the univariate FPS with non-negative real coefficients.
Definition 3.2 ($U(R)$ 1)
Let $R>0$ be a real number and $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$ be in $\mathbb{R}_{\geq 0}[[x]]$. We define
${\textstyle U(R)=\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}R^{j}=\sum_{n=0}^{\infty}u_{n}R^{n}\,.}$
Thus $U(R)$ is a nonnegative real number or +$\infty$+ ∞.
The limit always exists because partial sums $\sum_{j=0}^{n}u_{j}R^{j}$ are non-decreasing.
Proposition 3.3
Suppose that $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$ and $V(x)=\sum_{n=0}^{\infty}v_{n}x^{n}$ are in $\mathbb{R}_{\geq 0}[[x]]$ and that formally $U(x)=V(x)\frac{1}{1-x}$. Then
$\lim_{n\to\infty}u_{n}=\lim_{n\to\infty}(v_{0}+v_{1}+\dots+v_{n})=V(1)\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
Proof. This is immediate from the equality $u_{n}=\sum_{j=0}^{n}v_{j}$, $n\in\mathbb{N}_{0}$. $\Box$
For the sake of completeness we prove the next folklore result which we call the weak Abel theorem. The full Abel theorem [2] is Theorem 5.14.
Theorem 3.4 (weak Abel)
Let $R>0$ be a real number and let $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$ be in $\mathbb{R}_{\geq 0}[[x]]$. Suppose that for every $x\in[0,R)$ the series $U(x)$ converges and consider its sum $F_{U}\colon[0,R)\to[0,+\infty)$. Then
$\lim_{x\to R}F_{U}(x)=U(R)\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
Proof. Let $R$ and $U(x)$ be as stated and $N\in\mathbb{N}$. Since $R$ and the coefficients $u_{n}$ are nonnegative,
${\textstyle\sum_{n=0}^{N}u_{n}R^{n}=\lim_{x\to R^{-}}\sum_{n=0}^{N}u_{n}x^{n}\leq}$
${\textstyle\leq\lim_{x\to R}F_{U}(x)=\lim_{x\to R}\sum_{n=0}^{\infty}u_{n}x^{n}\leq\sum_{n=0}^{\infty}u_{n}R^{n}=U(R)\,.}$
For $N\to+\infty$ we get that $U(R)=\lim_{x\to R}F_{U}(x)$. $\Box$
We obtain two corollaries of the weak Abel theorem.
Corollary 3.5
Suppose that $U(x)$ and $V(x)$ in $\mathbb{R}_{\geq 0}[[x]]$ converge on $[0,1)$, $U(0)=u_{0}>0$ and that formally $V(x)=1-\frac{1}{U(x)}$. Then $U(1)>0$ and
${\textstyle V(1)=1-\frac{1}{U(1)}\,.}$
In particular, if $U(1)=+\infty$ then $V(1)=1$.
Proof. Formally $V(x)U(x)=U(x)-1$. It follows that $F_{U}>0$ on $[0,1)$ and that $F_{U}$ is non-decreasing. By Propositions 5.3, 5.5 and 5.6 (which we prove later),
$F_{V}(x)F_{U}(x)=F_{U}(x)-1$
and $F_{V}(x)=1-\frac{1}{F_{U}(x)}$ for every $x\in[0,1)$. For $x\to 1$ we get by Theorem 3.4 that $U(1)>0$ and $V(1)=1-\frac{1}{U(1)}$. In particular, if $U(1)=\lim_{x\to 1}F_{U}(x)=+\infty$ then $V(1)=\lim_{x\to 1}\big{(}1-\frac{1}{F_{U}(x)}\big{)}=1$. $\Box$
Corollary 3.6
Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{R}_{\geq 0}[[x]]$ and converge on $[0,1)$, $W(0)=w_{0}>0$, $W(1)<+\infty$ and that formally $W(x)=V(x)+W(x)U(x)^{2}$. Then $W(1)>0$, $\frac{V(1)}{W(1)}\leq 1$ and
${\textstyle U(1)=\sqrt{1-\frac{V(1)}{W(1)}}\,.}$
Proof. $W(1)>0$ is immediate from $w_{0}>0$. Arguing as in the previous proof we get the equality $W(1)=V(1)+W(1)U(1)^{2}$. Since $U(1)\geq 0$, we have $U(1)=\sqrt{1-\frac{V(1)}{W(1)}}$. $\Box$
An alternative derivation of Corollaries 3.5 and 3.6 avoids functions $F_{U}(x)$, $F_{V}(x)$, and $F_{W}(x)$ by only manipulating series with nonnegative summands, convergent ones and those with the sum +$\infty$+ ∞
We review the asymptotic symbols $\ll$, $\gg$, $O$ and $\sim$. Let $f,g\colon X\to\mathbb{C}$ where $X$ is a nonempty set. We say that $f\ll g$ on $X$ if there is a real constant $c>0$ such that $|f(x)|\leq c\cdot|g(x)|$ for every $x\in X$. The notation $f\gg g$ means that $g\ll f$, and $f=O(g)$ is synonymous with $f\ll g$. If $X\subset\mathbb{R}$ and $A\in\mathbb{R}\cup\{-\infty,+\infty\}$ is a limit point of the definition domain of $\frac{f}{g}$, we write $f\sim g$ ($x\to A$) if $\lim_{x\to A}\frac{f(x)}{g(x)}=1$. If $X=\mathbb{N}$, $A=+\infty$ and $f\sim g$ ($x\to+\infty$) then both $f\ll g$ and $f\gg g$ hold on the definition domain of $\frac{f}{g}$.
Propositions 3.7 and 3.8 below deal with the above generating function $B(x)$. We show that $B(1)=+\infty$ for $d\leq 2$ and $B(1)<+\infty$ for $d\geq 3$. These are well known results. The reason for proving them here in detail is that we want to have proofs for them that we like. Let $b_{n}$, $d_{n}$ and $B(x)$ be as above.
Proposition 3.7
Let $d\leq 2$. Then $b_{n}=0$ iff $n\in\mathbb{N}$ is odd, and
${\textstyle\frac{b_{n}}{d_{n}}=\frac{b_{n}}{(2d)^{n}}\gg n^{-d/2}}$
on even $n\in\mathbb{N}_{0}$. Hence if $d\leq 2$ then $B(1)=+\infty$.
Proof. The fact that the $b_{n}$ are supported on even $n$ holds in every dimension because every walk in $G_{d}$ that starts and ends at $\overline{0}$ does in each of the $d$ directions equally many back and forth steps. For even $n$ obviously $b_{n}\geq 1$. As for the estimate $\gg$, we prove the stronger relation $\sim$.
Let $n=2m$ with $m$ in $\mathbb{N}_{0}$ and $d=1$. Then $b_{n}=b_{2m}=\binom{2m}{m}$. Hence $\frac{b_{n}}{d_{n}}=4^{-m}\binom{2m}{m}$. The Stirling formula $m\sim\sqrt{2\pi m}\big{(}\frac{m}{\mathrm{e}}\big{)}^{m}$ ($m\to+\infty$), see [27] and Theorem 4.1, gives
${\textstyle\frac{b_{n}}{d_{n}}=4^{-m}\cdot(2m)!\cdot m!^{-2}\sim\frac{1}{\sqrt{\pi}}\cdot m^{-1/2}=\sqrt{2/\pi}\cdot n^{-1/2}\ \ (n\to+\infty)\,.}$
Let $d=2$. Then
${\textstyle b_{n}=b_{2m}=\sum_{j=0}^{m}\frac{(2m)!}{j!(m-j)!j!(m-j)!}=\binom{2m}{m}\sum_{j=0}^{m}\binom{m}{j}^{2}=\binom{2m}{m}^{2}\,.}$
The ratio is the number of walks in $G_{2}$ starting and ending at $\overline{0}$ and doing $j$ steps left, $j$ steps right, $m-j$ steps up and $m-j$ steps down. The last equality follows from the well-known binomial identity $\sum_{j=0}^{m}\binom{m}{j}^{2}=\binom{2m}{m}$. The asymptotic relation
${\textstyle\frac{b_{n}}{d_{n}}=\binom{2m}{m}^{2}\cdot 4^{-2m}\sim\frac{1}{\pi m}=\frac{2}{\pi}\cdot n^{-1}\ \ (n\to+\infty)}$
follows again from the Stirling formula. By Definition 3.2 and since $\sum_{n=1}^{\infty}n^{-1}=+\infty$, for $d\leq 2$ we have $B(1)=+\infty$. $\Box$
Proposition 3.8
If $d\geq 3$ then
${\textstyle\frac{b_{n}}{d_{n}}\ll n^{-d/2}}$
on $\mathbb{N}$. Hence for $d\geq 3$ we have $B(1)<+\infty$.
Proof. Let $d\geq 3$. As we know, $b_{n}=0$ for odd $n$. For $n=2m$ with $m\in\mathbb{N}_{0}$ the fraction $\frac{b_{n}}{d_{n}}=\frac{b_{2m}}{d_{2m}}$ equals
${\textstyle(2d)^{-2m}\sum_{\begin{subarray}{c}m_{i}\in\mathbb{N}_{0}\\ m_{1}+\dots+m_{d}=m\end{subarray}}\frac{(2m)!}{(m_{1}!\cdot\ldots\cdot m_{d}!)^{2}}=4^{-m}\binom{2m}{m}\sum_{\cdots}\big{[}\frac{1}{d^{m}}\binom{m}{m_{1},\,\dots,\,m_{d}}\big{]}^{2}}$
$\leq$ ${\textstyle 4^{-m}\binom{2m}{m}\cdot\max_{\cdots}d^{-m}\binom{m}{m_{1},\,\dots,\,m_{d}}=4^{-m}\binom{2m}{m}\cdot d^{-m}\binom{m}{m_{1}^{\prime},\,\dots,\,m_{d}^{\prime}}}$
$=$ ${\textstyle\frac{1}{\sqrt{\pi m}}\cdot\mathrm{e}^{O(1)}\cdot m^{(1-d)/2}\ll n^{-d/2},\ \text{$m_{i}^{\prime}\in\{\lfloor\frac{m}{d}\rfloor,\lceil\frac{m}{d}\rceil\}$ and $\sum_{i=1}^{d}m_{i}^{\prime}=m$}\,.}$
The numbers of $\lfloor\frac{m}{d}\rfloor$s and $\lceil\frac{m}{d}\rceil$s in the tuple $m_{1}^{\prime},\dots,m_{d}^{\prime}$ are uniquely determined. In the first line we count the walks as in the case $d\leq 2$. The numbers in []delimited-[][$\dots$][ … ] sum to $1$ due to the multinomial expansion of $d^{m}=(1+\dots+1)^{m}$. The inequality in the second line follows from the lemma that if $a_{1},\dots,a_{r}\geq 0$ and $a_{1}+\dots+a_{r}=1$ then $a_{1}^{2}+\dots+a_{r}^{2}\leq\max_{1\leq i\leq r}a_{i}$. The multinomial coefficient is maximized by the lemma that if $a>b\geq 0$ are integers with $a\geq b+2$ then $a!\cdot b!>(a-1)!\cdot(b+1)!$. In the third line we apply in the following way the asymptotics from the proof of Proposition 3.7 and the Stirling formula. As $m\to+\infty$,
${\textstyle\binom{m}{m_{1}^{\prime},\,\dots,\,m_{d}^{\prime}}=\frac{m!}{\prod_{i=1}^{d}m_{i}^{\prime}!}\sim\frac{\sqrt{2\pi m}\cdot(m/\mathrm{e})^{m}}{\prod_{i=1}^{d}\sqrt{2\pi m_{i}^{\prime}}\cdot(m_{i}^{\prime}/\mathrm{e})^{m_{i}^{\prime}}}=\mathrm{e}^{O(1)}\cdot m^{(1-d)/2}\cdot d^{m}\,,}$
because $\sum_{i=1}^{d}m_{i}^{\prime}=m$, $\sqrt{m_{i}^{\prime}}=(1+O(m^{-1}))\sqrt{m/d}$, $\log(m_{i}^{\prime})=\log(\frac{m}{d})+O(m^{-1})$ and
${\textstyle\big{(}m_{i}^{\prime}\big{)}^{m_{i}^{\prime}}=\mathrm{e}^{m_{i}^{\prime}\log(m_{i}^{\prime})}=\big{(}\frac{m}{d}\big{)}^{m_{i}^{\prime}}\cdot\mathrm{e}^{O(m_{i}^{\prime}/m)}\,.}$
Here $\mathrm{e}^{O(1)}$ denotes a function $f\colon\mathbb{N}_{0}\to(0,+\infty)$ such that $1\ll f\ll 1$ on $\mathbb{N}_{0}$. Since the series $\sum_{n=1}^{\infty}n^{-3/2}$ absolutely converges, for $d\geq 3$ we have $B(1)<+\infty$. $\Box$
We obtain a main result of our article, a strengthening of Theorem 1.1 on $\overline{0}$-recurrence.
Theorem 3.9 ($\overline{0}$-recurrence+)
Let $d\in\mathbb{N}$. Using the above notation we have
$\lim_{n\to\infty}\frac{a_{n}}{d_{n}}=\lim_{n\to\infty}\frac{|W_{d,\mathrm{rec}}(n)|}{(2d)^{n}}\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ =1-\frac{1}{B(1)}\in(0,\,1)&\dots&\text{if $d\geq 3$}\,,\end{array}\right.$
where $B(x)$ is the above generating function of closed walks.
Proof. By part 1 of Proposition 3.1, formally $A(x)=C(x)\frac{1}{1-1x}$. By Proposition 3.3, $\lim_{n\to\infty}\frac{a_{n}}{d_{n}}=C(1)$. By part 2 of Proposition 3.1 and Corollary 3.5, $C(1)=1-\frac{1}{B(1)}$. If $d\leq 2$ then by Proposition 3.7 and Corollary 3.5, $C(1)=1$. If $d\geq 3$ then by Proposition 3.8 and Corollary 3.5, $C(1)=1-\frac{1}{B(1)}\in(0,1)$. $\Box$
Next we strengthen Theorem 1.2 on $\overline{v}$-recurrence, the theorem G. Pólya proved in [23]. Let $d\in\mathbb{N}$, $\overline{v}\in\mathbb{Z}^{d}\setminus\{\overline{0}\}$ (we have just treated the case $\overline{v}=\overline{0}$) and $n\in\mathbb{N}_{0}$. We again define numbers of certain $\overline{0}$-walks of length $n$ ($\in\mathbb{N}_{0}$) in $G_{d}$. Let $a_{n}^{\prime}$ be the number of $\overline{v}$-recurrent walks, $b_{n}$ be as before the number of walks ending at $\overline{0}$, $b_{n}^{\prime}$ be the number of walks ending at $\overline{0}$ and avoiding $\overline{v}$, $c_{n}^{\prime}$ be the number of walks ending at $\overline{v}$ and with the inner avoiding $\overline{v}$, $c_{n}^{\prime\prime}$ be the number of walks ending at $\overline{v}$ and with the inner avoiding both $\overline{0}$ and $\overline{v}$, and $d_{n}=(2d)^{n}$ be as before the number of all walks. We have $a_{0}^{\prime}=c_{0}^{\prime}=c_{0}^{\prime\prime}=0$ and $b_{0}=b_{0}^{\prime}=d_{0}=1$. Besides the previous generating functions $B(x)$ and $D(x)=\frac{1}{1-x}$ we consider the generating functions
${\textstyle A_{0}(x)=\sum_{n=0}^{\infty}\frac{a_{n}^{\prime}}{d_{n}}x^{n},\ B_{0}(x)=\sum_{n=0}^{\infty}\frac{b_{n}^{\prime}}{d_{n}}x^{n},\ C_{0}(x)=\sum_{n=0}^{\infty}\frac{c_{n}^{\prime}}{d_{n}}x^{n}}$
and ${\textstyle\text{and }C_{1}(x)=\sum_{n=0}^{\infty}\frac{c_{n}^{\prime\prime}}{d_{n}}x^{n}\,.}$
The dependence on $d$ and $\overline{v}$ is left implicit in the notation. The quantities we want to determine are $a_{n}^{\prime}$ and $A_{0}(x)$.
We obtain an analog of Proposition 3.1.
Proposition 3.10
Let $A_{0}(x)$, $B(x)$, $B_{0}(x)$, $C_{0}(x)$, $C_{1}(x)$, and $D(x)$ be the above generating functions. The following formal relations hold.
  1. 1.
    $A_{0}(x)=C_{0}(x)D(x)=C_{0}(x)\frac{1}{1-x}$.
  2. 2.
    $B(x)=B_{0}(x)+C_{0}(x)^{2}B(x)$, equivalently $B(x)=\frac{B_{0}(x)}{1-C_{0}(x)^{2}}$.
  3. 3.
    $C_{0}(x)=B_{0}(x)\cdot C_{1}(x)$.
Proof. 1. The walks $w$ counted by $A_{0}(x)$ split uniquely at the first visit of $\overline{v}$ as $w=w_{1}w_{2}$, where $w_{1}$ are walks counted by $C_{0}(x)$ and $w_{2}$ are arbitrary $\overline{v}$-walks. By the vertex-transitivity of $G_{d}$, the walks $w_{2}$ are counted by $D(x)$.
2. It suffices to prove the first equality. The generating function $B(x)$ counts $\overline{0}$-walks $w$ in $G_{d}$ ending at $\overline{0}$. Those avoiding $\overline{v}$ are counted by $B_{0}(x)$, and those visiting $\overline{v}$ uniquely decompose as $w=w_{1}w_{2}w_{3}$, where the walk $w_{1}$ starts at $\overline{0}$, ends at $\overline{v}$ and the inner avoids $\overline{v}$, the walk $w_{2}$ starts and ends at $\overline{v}$, and the walk $w_{3}$ starts at $\overline{v}$, ends at $\overline{0}$ and the inner avoids $\overline{v}$. Reversing $w_{3}$ we see that both $w_{1}$ and $w_{3}$ are counted by $C_{0}(x)$. By the vertex-transitivity of $G_{d}$, the middle walks $w_{2}$ are counted by $B(x)$. We get the first equality.
3. We consider the walks $w$ counted by $C_{0}(x)$. They start at $\overline{0}$, end at $\overline{v}$ and the inner avoids $\overline{v}$. Each $w$ uniquely splits at the last visit of $\overline{0}$ in two walks as $w=w_{1}w_{2}$, where $w_{1}$ and $w_{2}$ are counted by $B_{0}(x)$ and $C_{1}(x)$, respectively. The relation follows. $\Box$
We obtain a main result of our article, a strengthening of Theorem 1.2 on $\overline{v}$-recurrence.
Theorem 3.11 ($\overline{v}$-recurrence+)
Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$, $\overline{v}\neq\overline{0}$. Using the above notation we have
$\lim_{n\to\infty}\frac{a_{n}^{\prime}}{d_{n}}=\lim_{n\to\infty}\frac{|W_{d}(\overline{v},\,n)|}{(2d)^{n}}\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ =\sqrt{1-\frac{B_{0}(1)}{B(1)}}\in(0,\,1)&\dots&\text{if $d\geq 3$}\,,\end{array}\right.$
where $B_{0}(x)$ and $B(x)$ are the above generating functions of closed walks.
Proof. The generating functions $B(x)$, $B_{0}(x)$, $C_{0}(x)$ and $C_{1}(x)$ have all coefficients in $[0,1]$. Thus their sums are continuous and non-decreasing functions
[…]
We show that $F_{C_{0}}$ is bounded on $[0,1)$. If $C_{0}(x)=0$ then it certainly holds. Suppose that $C_{0}(x)$ has at least one positive coefficient (in fact, it has infinitely many of them, but we give our argument with minimal assumptions so that we can use it later in more general situations). If $F_{C_{0}}<1$ on $[0,1)$, we are done. Else, since $F_{C_{0}}(0)=0$ and $F_{C_{0}}$ is continuous and increasing, there is a point $y\in(0,1)$ such that $F_{C_{0}}<1$ on $[0,y)$ and $\lim_{x\to y^{-}}F_{C_{0}}(x)=1$. But then, since $F_{B_{0}}(0)=1$, part 2 of Proposition 3.10 gives
$\lim_{x\to y^{-}}F_{B}(x)=+\infty$
and this contradicts the continuity of $F_{B}$ at $y$. Hence $F_{C_{0}}$ is bounded on $[0,1)$. It follows that $F_{B_{0}}$ is bounded on $[0,1)$ — else part 3 of Proposition 3.10 gives, since $C_{1}(x)$ has at least one positive coefficient, that
$\lim_{x\to 1}F_{C_{0}}(x)=\lim_{x\to 1}F_{B_{0}}(x)\cdot\lim_{x\to 1}F_{C_{1}}(x)=+\infty\,.$
Let $d\leq 2$. Taking the limit for $x\to 1$ we get, due to the boundedness of $F_{B_{0}}$, Proposition 3.7, part 2 of Proposition 3.10 and Theorem 3.4, that $C_{0}(1)=1$. By part 1 of Proposition 3.10 and Proposition 3.3,
${\textstyle\lim_{n\to\infty}\frac{a_{n}^{\prime}}{d_{n}}=C_{0}(1)=1\,.}$
Let $d\geq 3$. Part 2 of Proposition 3.10, Proposition 3.8 and Corollary 3.6 give $C_{0}(1)=\sqrt{1-\frac{B_{0}(1)}{B(1)}}$. By part 1 of Proposition 3.10 and Proposition 3.3 we have
${\textstyle\lim_{n\to\infty}\frac{a_{n}^{\prime}}{d_{n}}=C_{0}(1)=\sqrt{1-\frac{B_{0}(1)}{B(1)}}\,.}$
Clearly, $0<B_{0}(1)<B(1)<+\infty$. Thus the root lies in $(0,1)$. $\Box$
Corollary 3.12
If $d\leq 2$ then $\lim_{x\to 1}F_{B}(x)=+\infty$ and $\lim_{x\to 1}\frac{F_{B_{0}}(x)}{F_{B}(x)}=0$.
Proof. Let $d\leq 2$. The first limit follows from Theorem 3.4 and Proposition 3.7. Part 2 of Proposition 3.10, and Propositions 5.3, 5.5 and 5.6 imply that for every $x\in[0,1)$,
${\textstyle\frac{F_{B_{0}}(x)}{F_{B}(x)}=}1-F_{C_{0}}(x)^{2}\,.$
In the previous proof we showed that for $d\leq 2$, $\lim_{x\to 1}F_{C_{0}}(x)=1$. The stated limit follows. $\Box$
If $d\geq 3$ and one sets $\overline{v}=\overline{0}$, then Theorem 3.11 does not turn in Theorem 3.9: $B_{0}(1)$ turns in $C(1)$ and, by part 2 of Proposition 3.1,
${\textstyle\sqrt{1-\frac{C(1)}{B(1)}}=\sqrt{\big{(}1-\frac{1}{B(1)}\big{)}^{2}+\frac{1}{B(1)}}>1-\frac{1}{B(1)}\,.}$
4 Effective Theorem 1.1 for $d=2$
Let $f\colon\mathbb{N}\to\mathbb{R}$. The bare limit
$\lim_{n\to\infty}f(n)=c$
is an ineffective result, it does not tell us anything about the distance between $f(n)$ and $c$ for any concrete $n\in\mathbb{N}$. Theorems 1.1, 1.2, 3.9 and 3.11 on $\overline{0}$- and $\overline{v}$-recurrence are in this ineffective form. We improve upon it a little and in Propositions 4.3 and 4.5 we effectivize Theorem 1.1 if $d=2$. We make use of an effective Stirling formula.
Theorem 4.1 (Robbins [27])
For every $n\in\mathbb{N}$,
${\textstyle\sqrt{2\pi n}\cdot\big{(}\frac{n}{\mathrm{e}}\big{)}^{n}\cdot\mathrm{e}^{1/(1+12n)}\leq n!\leq\sqrt{2\pi n}\cdot\big{(}\frac{n}{\mathrm{e}}\big{)}^{n}\cdot\mathrm{e}^{1/12n}\,.}$
For example,
$5039.33\ldots\leq 7!=5040\leq 5040.04\dots\,.$
We obtain an effective version of Corollary 3.5 in the case $U(1)=+\infty$.
Proposition 4.2
Suppose that the FPS
${\textstyle U(x)=\sum_{n\geq 0}u_{n}x^{n}\,\text{ and }\,V(x)=\sum_{n\geq 0}v_{n}x^{n}}$
in $\mathbb{R}_{\geq 0}[[x]]$ absolutely converge for every $x\in[0,1)$. Let $u_{0}>0$, $v_{n}\leq 1$ and $u_{2n}\geq\frac{c}{n}$ for every $n$ in $\mathbb{N}$ and some constant $c>0$, and let formally $V(x)=1-\frac{1}{U(x)}$. Then $V(1)=\sum_{n=0}^{\infty}v_{n}=1$ and for every $N\in\mathbb{N}$ and $x\in(\frac{1}{2},1)$ we have
$1\geq\sum_{n=0}^{N}v_{n}\geq 1-\frac{1}{c\log\big{(}\frac{1}{1-x}\big{)}-c\log 2}-\frac{x^{N+1}}{1-x}\,.$
Proof. Since, by the assumption, $U(1)=+\infty$, Corollary 3.5 yields $V(1)=1$. This gives the first displayed inequality because $u_{n}\geq 0$. We prove the second one. Since $u_{2n}\geq\frac{c}{n}$, for every $x\in(0,1)$ we have
${\textstyle F_{U}(x)=\sum_{n=0}^{\infty}u_{n}x^{n}\geq c\sum_{n=1}^{\infty}\frac{x^{2n}}{n}=c\log\big{(}\frac{1}{1-x^{2}}\big{)}\,.}$
Since $v_{n}\leq 1$ and for every $x\in[0,1)$ we have (by the proof of Corollary 3.5) $F_{V}(x)=1-\frac{1}{F_{U}(x)}$, for every $N\in\mathbb{N}$ and $x\in(\frac{1}{2},1)$ we have
${\textstyle\sum_{n=0}^{N}v_{n}\geq\sum_{n=0}^{N}v_{n}x^{n}=\sum_{n=0}^{\infty}v_{n}x^{n}-\sum_{n>N}v_{n}x^{n}}$
${\textstyle\geq 1-\frac{1}{F_{U}(x)}-\frac{x^{N+1}}{1-x}\geq 1+\frac{1}{c\log(1-x)+c\log(1+x)}-\frac{x^{N+1}}{1-x}}$
${\textstyle\geq 1-\frac{1}{c\log(\frac{1}{1-x})-c\log 2}-\frac{x^{N+1}}{1-x}\,.}$
$\Box$
Thus we obtain an effective upper bound on the difference
$1-4^{-N}|W_{\mathrm{2,rec}}(N)|$
in Theorem 1.1 for $d=2$.
Proposition 4.3
For every $N\geq 3$ we have
$0\leq 1-4^{-N}|W_{\mathrm{2,rec}}(N)|\leq(0.2\log N-0.16)^{-1}+N^{8/9}\exp(-N^{1/9})\;.$
Proof. Let $d=2$, $n\in\mathbb{N}_{0}$ and let $b_{n}$ and $c_{n}$ be as defined at the beginning of Section 3. We set $v_{n}=c_{n}4^{-n}$ and $u_{n}=b_{n}4^{-n}$. Then $v_{n},u_{n}\geq 0$, $v_{n}\leq 1$, $u_{0}>0$, the power series $V(x)=C(x)$ and $U(x)=B(x)$ absolutely converge on $[0,1)$, and we have the formal relation $V(x)=1-\frac{1}{U(x)}$. We bound $u_{2n}$ from below by Theorem 4.1: for every $n\in\mathbb{N}$,
${\textstyle u_{2n}=4^{-2n}\frac{(2n)!^{2}}{n!^{4}}\geq\frac{\mathrm{e}^{2/(24n+1)}}{\pi n\mathrm{e}^{1/3n}}\geq\frac{1}{\pi n\mathrm{e}^{1/3}}}$
and $u_{2n}\geq\frac{0.228}{n}$. For $N\in\mathbb{N}$ we set $x=x(N)=1-N^{-8/9}$ ($\in[0,1)$). Then for $N\geq 2$,
${\textstyle\frac{x^{N+1}}{1-x}}$ $=$ ${\textstyle\frac{N^{8/9}}{\exp((N+1)\log(1/(1-N^{-8/9})))}}$
$=$ ${\textstyle\frac{N^{8/9}}{\exp((N+1)(N^{-8/9}+N^{-2\cdot 8/9}/2+\dots))}\leq\frac{N^{8/9}}{\exp(N^{1/9})}}\,.$
For $N\geq 3$ we have $x\in(\frac{1}{2},1)$ and $\log(\frac{1}{1-x})=\frac{8}{9}\log N$. We use Proposition 4.2 with $c=0.228$. Then $\frac{8}{9}c\geq 0.2$ and $c\log 2\leq 0.16$, and the equality $\frac{a_{n}}{4^{n}}=\sum_{n=0}^{N}\frac{c_{n}}{4^{n}}$ in the proof of part 1 of Proposition 3.1 gives the stated bound that for every integer $N\geq 3$,
${\textstyle 1\geq 4^{-N}|W_{\mathrm{2,rec}}(N)|=\sum_{n=0}^{N}v_{n}\geq 1-\frac{1}{0.2\log N-0.16}-\frac{N^{8/9}}{\exp(N^{1/9})}\,.}$
$\Box$
We adapt Proposition 4.2 for the bound $u_{2n}\leq\frac{c^{\prime}}{n}$.
Proposition 4.4
Suppose that the FPS
${\textstyle U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}\,\text{ and }\,V(x)=\sum_{n=0}^{\infty}v_{n}x^{n}}$
in $\mathbb{R}_{\geq 0}[[x]]$ absolutely converge for every $x\in[0,1)$. Let $u_{0}=1$, $u_{2n-1}=0$ and $u_{2n}\leq\frac{c^{\prime}}{n}$ for every $n\in\mathbb{N}$ and some constant $c^{\prime}>0$, and let formally $V(x)=1-\frac{1}{U(x)}$. Then for every $N\in\mathbb{N}$ and $x\in(0,1)$ with $2N(1-x)\leq\frac{1}{3}$ we have
$\sum_{n=0}^{N}v_{n}\leq 1+3N(1-x)-\frac{1}{\big{(}c^{\prime}+\frac{1}{\log(6N)}\big{)}\cdot\log\big{(}\frac{1}{1-x}\big{)}}\;.$
Proof. By the assumptions $1\leq\frac{\log(1/(1-x))}{\log(6N)}$ and $x^{2n}\leq x^{n}$. So for every $x$ as stated we have
${\textstyle F_{U}(x)=\sum_{n=0}^{\infty}u_{n}x^{n}\leq 1+c^{\prime}\sum_{n=1}^{\infty}\frac{x^{2n}}{n}\leq\big{(}c^{\prime}+\frac{1}{\log(6N)}\big{)}\log\big{(}\frac{1}{1-x}\big{)}\,.}$
For every $N$ in $\mathbb{N}$ and $x=1-y\in(0,1)$ with $y\in(0,1)$ close to 00 so that $2Ny\leq\frac{1}{3}$ we therefore have, since $u_{n}\geq 0$, the stated bound
${\textstyle\sum_{n=0}^{N}v_{n}}$$\leq$ ${\textstyle x^{-N}\sum_{n=0}^{\infty}v_{n}x^{n}=\frac{1-1/F_{U}(x)}{x^{N}}\leq\frac{1}{x^{N}}\big{(}1+\frac{1}{(c^{\prime}+\frac{1}{\log(6N)})\log(1-x)}\big{)}}$
$=$ $\mathrm{e}^{N\log(1/(1-y))}(\cdots)\leq\mathrm{e}^{2Ny}(\cdots)\leq(1+3Ny)(\cdots)$
$\leq$ ${\textstyle 1+3N(1-x)-\frac{1}{(c^{\prime}+\frac{1}{\log(6N)})\log(\frac{1}{1-x})}\,.}$
$\Box$
Thus we obtain an effective lower bound on the difference
$1-4^{-N}|W_{\mathrm{2,rec}}(N)|$
in Theorem 1.1 for $d=2$.
Proposition 4.5
If $N\geq 140000$ then
$1-4^{-N}|W_{\mathrm{2,rec}}(N)|\geq(0.84\log N)^{-1}-3N^{-1}\;.$
Proof. Again $v_{n}=c_{n}4^{-n}$ and $u_{n}=b_{n}4^{-n}$, so that $V(x)=C(x)$ and $U(x)=B(x)$. The assumptions of the previous proposition are satisfied if we get the required upper bound on $u_{2n}$. By Theorem 4.1 for every $n\in\mathbb{N}$,
${\textstyle u_{2n}=4^{-2n}\frac{(2n)!^{2}}{n!^{4}}\leq\frac{\mathrm{e}^{1/12n}}{\pi n\mathrm{e}^{4/(12n+1)}}\leq\frac{\mathrm{e}^{1/12}}{\pi n}}$
and $u_{2n}\leq\frac{0.346}{n}$. We set $x=x(N)=1-N^{-2}\in[0,1)$. For $N\geq 1.4\cdot 10^{5}$ one has $2N(1-x)\leq\frac{1}{3}$ and $\frac{1}{\log(6N)}\leq 0.074$. For $c^{\prime}=0.346$ the previous proposition gives for $4^{-N}a_{N}=\sum_{n=0}^{N}v_{n}$ the stated bound. $\Box$
We combine both bounds in one estimate.
Corollary 4.6
Let $d=2$. There is an effective $N_{0}$ such that for every $N\geq N_{0}$,
$(0.9\log N)^{-1}\leq 1-4^{-N}|W_{\mathrm{2,rec}}(N)|\leq(0.1\log N)^{-1}\,.$
5 The model $\mathrm{W_{\mathbb{C}}}$
Working in the model $\mathrm{W_{\mathbb{C}}}$ we generalize Theorem 3.9 on $\overline{0}$-recurrence to Theorems 5.18 and 5.19, and Theorem 3.11 on $\overline{v}$-recurrence to Theorems 5.22 and 5.23. All four theorems go beyond probability because they work with general weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We use absolute convergence of complex series introduced in Section 1.
We begin with an infinite triangle inequality and then introduce sum, linear combination, grouping and product of complex series. The proofs for the first two operations are omitted.
Proposition 5.1
Let $\sum_{x\in X}h(x)$ be an absolutely convergent series. Then we have the inequality for sums
${\textstyle\big{|}\sum_{x\in X}h(x)\big{|}\leq\sum_{x\in X}|h(x)|\,.}$
Proof. Let $\sum_{x\in X}h(x)$ be as stated. For finite $X$ the inequality is the standard triangle inequality in $\mathbb{C}$. Suppose that $X$ is countable and that the series has the sum $s$ ($\in\mathbb{C}$). Then also $\sum_{x\in X}|h(x)|$ absolutely converges and has the sum $t$ ($\in\mathbb{R}_{\geq 0}$). For any bijection $f\colon\mathbb{N}\to X$,
$|s|$ $=$ $\big{|}\lim_{n\to\infty}{\textstyle\sum_{j=1}^{n}h(f(j))}\big{|}=\lim_{n\to\infty}{\textstyle\big{|}\sum_{j=1}^{n}h(f(j))\big{|}}$
$\leq$ $\lim_{n\to\infty}{\textstyle\sum_{j=1}^{n}|h(f(j))|=t\,,}$
due to continuity of the function $u\mapsto|u|$, $u\in\mathbb{C}$. $\Box$
If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are two series on two disjoint sets $X_{i}$, then the series $S_{1}+S_{2}=\sum_{x\in X_{1}\cup X_{2}}h(x)$, where $h(x)=h_{i}(x)$ if $x\in X_{i}$, is the sum of $S_{1}$ and $S_{2}$.
Proposition 5.2
If $S_{1}$ and $S_{2}$ absolutely converge and have respective sums $s_{1}$ and $s_{2}$, then $S_{1}+S_{2}$ absolutely converges and has the sum $s_{1}+s_{2}$.
If $\alpha,\beta\in\mathbb{C}$ and $S_{i}=\sum_{x\in X}h_{i}(x)$, $i=1,2$, are two series on a set $X$, then the series
${\textstyle\alpha S_{1}+\beta S_{2}=\sum_{x\in X}(\alpha h_{1}(x)+\beta h_{2}(x))}$
is the linear combination of $S_{1}$ and $S_{2}$.
Proposition 5.3
If $S_{1}$ and $S_{2}$ absolutely converge and have respective sums $s_{1}$ and $s_{2}$, then $\alpha S_{1}+\beta S_{2}$ absolutely converges and has the sum $\alpha s_{1}+\beta s_{2}$.
For the remaining two operations we need a lemma.
Lemma 5.4
Let $S=\sum_{x\in X}h(x)$ be an absolutely convergent series with the sum $s$. Then for every $\varepsilon>0$ there is a finite set $Y\subset X$, denoted by $Y(S,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
${\textstyle\big{|}s-\sum_{x\in Z}h(x)\big{|}\leq\varepsilon\,.}$
Proof. For finite $X$ we take $Y=X$. For countable $X$ we take any bijection $f\colon\mathbb{N}\to X$. For the given $\varepsilon>0$ we take an $N\in\mathbb{N}$ such that $|\sum_{i=1}^{N}h(f(i))-s|\leq\frac{\varepsilon}{2}$ and $\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}$. We set $Y=f[\,[N]\,]$. Then for every finite set $Z$ with $Y\subset Z\subset X$,
${\textstyle|\sum_{x\in Z}h(x)-s|\leq|\sum_{x\in Y}h(x)-s|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.}$
$\Box$
partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ of $S$ absolutely converges and has the sum $s_{Z}$, then the series
${\textstyle S_{P}=\sum_{Z\in P}s_{Z}=\sum_{Z\in P}\sum_{x\in Z}h(x)}$
(the inner $\sum$ means sum) is the $P$-grouping of $S$.
Proposition 5.5
Suppose that the series $S=\sum_{x\in X}h(x)$ absolutely converges, has the sum $s$ and that $P$ is a partition of $X$. Then $S_{P}$ is correctly defined, absolutely converges, and has the same sum $s$ as $S$.
Proof. Let $S$, $s$ and $P$ be as stated. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined.
We show that $S_{P}$ absolutely converges. Let
${\textstyle c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and is finite}\})}$
and let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. Let $i\in[n]$. We use Lemma 5.4 and take the finite set $Z_{i}^{\prime}=Y(S_{Z_{i}},2^{-i})$ ($\subset Z_{i}$). We set $Z_{0}=Z_{1}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$); this is a disjoint union. Then, in the sense of sums,
${\textstyle\sum_{i=1}^{n}\big{|}\sum_{x\in Z_{i}}h(x)\big{|}}$$\leq$ ${\textstyle\sum_{i=1}^{n}\big{|}\sum_{x\in Z_{i}}h(x)-\sum_{x\in Z_{i}^{\prime}}h(x)\big{|}\,+}$
$+$ ${\textstyle\sum_{x\in Z_{0}}|h(x)|\leq\sum_{i=1}^{n}2^{-i}+c}$
$\leq$ $1+c\,.$
Thus $S_{P}$ absolutely converges.
Let $t$ be the sum of $S_{P}$. We show that $t=s$. It suffices to show that for any given $\varepsilon$ we have $|t-s|\leq\varepsilon$. We use Lemma 5.4 and take finite sets $X^{\prime}\equiv Y(S,\frac{\varepsilon}{3})$ ($\subset X$) and $P^{\prime}\equiv Y(S_{P},\frac{\varepsilon}{3})$ ($\subset P$). Then we take a finite superset of blocks $\{Z_{1},\dots,Z_{n}\}$ of $P^{\prime}$, that is $P^{\prime}\subset\{Z_{1},\dots,Z_{n}\}\subset P$ and $n\in\mathbb{N}$, such that also $X^{\prime}\subset\bigcup_{i=1}^{n}Z_{i}$. We use Lemma 5.4 for every $i\in[n]$ and take the finite set $Z_{i}^{\prime}=Y(S_{Z_{i}},2^{-i}\frac{\varepsilon}{3})$ ($\subset Z_{i}$). Finally, for every $i\in[n]$ we set $Z_{i}^{\prime\prime}=Z_{i}^{\prime}\cup(X^{\prime}\cap Z_{i})$ ($\subset Z_{i}$) and take the disjoint union $X_{0}=\bigcup_{i=1}^{n}Z_{i}^{\prime\prime}$ ($\subset X$). Then $X_{0}$ is finite, $X^{\prime}\subset X_{0}$, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ and
$|s-t|$$\leq$ ${\textstyle|s-\sum_{x\in X_{0}}h(x)|+\sum_{i=1}^{n}|\sum_{x\in Z_{i}^{\prime\prime}}h(x)-\sum_{x\in Z_{i}}h(x)|\,+}$
$+$ ${\textstyle|\sum_{i=1}^{n}\sum_{x\in Z_{i}}h(x)-t|\leq\frac{\varepsilon}{3}+\frac{\varepsilon}{3}+\frac{\varepsilon}{3}=\varepsilon\,.}$
Hence $s=t$. $\Box$
If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are two series on two sets $X_{i}$, then the series
${\textstyle S_{1}\cdot S_{2}=\sum_{(x,y)\in X_{1}\times X_{2}}h_{1}(x)h_{2}(y)}$
is the product of $S_{1}$ and $S_{2}$.
Proposition 5.6
If $S_{1}$ and $S_{2}$ absolutely converge and have respective sums $s_{1}$ and $s_{2}$, then $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Suppose that $c\geq 0$ is such that for $i=1,2$ and any finite sets $Y_{i}$ with $Y_{i}\subset X_{i}$ one has $\sum_{x\in Y_{i}}|h_{i}(x)|\leq c$. Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
${\textstyle\sum_{(x,\,y)\in A}|h_{1}(x)h_{2}(y)|\leq\sum_{x\in Y_{1}}|h_{1}(x)|\cdot\sum_{y\in Y_{2}}|h_{2}(y)|\leq c\cdot c=c^{2}\,.}$
We show that the sum $s$ of $S_{1}\cdot S_{1}$ equals $s_{1}s_{2}$. It suffices to prove that for any given $\varepsilon\in(0,1]$ we have $|s-s_{1}s_{2}|\leq\varepsilon$. We use Lemma 5.4 and take finite sets $X^{\prime}_{1}=Y(S_{1},\frac{\varepsilon}{4(1+|s_{2}|)})$ ($\subset X_{1}$), $X^{\prime}_{2}=Y(S_{2},\frac{\varepsilon}{4(1+|s_{1}|)})$ ($\subset X_{2}$) and $Z=Y(S_{1}\cdot S_{2},\frac{\varepsilon}{4})$ ($\subset X_{1}\times X_{2}$). We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$. Then $|s-s_{1}s_{2}|$ is at most
${\textstyle\big{|}s-\sum_{(x,y)\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big{|}+\big{|}\sum_{x\in X_{1}^{\prime\prime}}h_{1}(x)\sum_{y\in X_{2}^{\prime\prime}}h_{2}(y)-s_{1}s_{2}\big{|}}$
${\textstyle\leq\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|,\ |t_{1}|\leq\frac{\varepsilon}{4(1+|s_{2}|)}\,\text{ and }\,|t_{2}|\leq\frac{\varepsilon}{4(1+|s_{1}|)}\,.}$
Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
As we know, $W(n)$ denotes the set of $1$-walks of length $n$ ($\in\mathbb{N}_{0}$) in $K_{\mathbb{N}}$. These are the tuples $w=\langle u_{0},u_{1},\dots,u_{n}\rangle$, $u_{i}\in\mathbb{N}$, such that $u_{0}=1$ and $u_{i-1}\neq u_{i}$ for every $i\in[n]$. Recall that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light if for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges.
We need sufficient conditions for lightness stated in terms of values of the given weight. (We make use of them later, in the follow-up article [16].) One such condition is that the graph $G(h)$ in Definition 1.3 has finite vertex degrees. We generalize it to a condition which can be used for graphs $G(h)$ with infinite degrees. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. By $V(h,n)$ ($\subset V(h)$) we denote the set of vertices in $G(h)$ lying in distance at most $n$ from the vertex $1$.
Definition 5.7 (slimness)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n\in\mathbb{N}_{0}$ there is a constant $c=c(n)\geq 0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
${\textstyle\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.}$
If every vertex in $G(h)$ has finite degree then $h$ is obviously slim.
Proposition 5.8
Every slim weight is light.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a slim weight. We show by induction on $n\in\mathbb{N}_{0}$ that for every $n$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges. For $n=0$ this is trivial. Let $n>0$ and let $c\geq 0$ be a constant such that for every finite set $X\subset W(n-1)$ we have $\sum_{w\in X}|h(w)|\leq c$ and that for every vertex $u\in V(h,n-1)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have $\sum_{v\in X}|h(\{u,\,v\})|\leq c$. Let $X\subset W(n)$ be a finite set. We may assume that every edge in every walk $w\in X$ has nonzero weight. Recall that $\ell(w)$ is the last vertex of a walk $w$. We decompose every walk $w\in X$ as $w=w^{\prime}\ell(w)$ with $w^{\prime}\in W(n-1)$. We denote the set of the walks $w^{\prime}$ by $Y$ and for any $w^{\prime}\in Y$ we set $X(w^{\prime})=\{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}:\;w^{\prime}v\in X\}$. Then the sum $\sum_{w\in X}|h(w)|$ is indeed at most
${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot\sum_{v\in X(w^{\prime})}|h(\{\ell(w^{\prime}),\,v\})|\leq\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c^{2}\,.}$
$\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a weight and $W$ be a set of walks in $K_{\mathbb{N}}$. If the series $\sum_{w\in W}h(w)$ absolutely converges, we denote its sum by $h(W)$ ($\in\mathbb{C}$). For $W=\emptyset$ we set $h(W)=0$. To extend Theorem 3.9 on $\overline{0}$-recurrence to $\mathrm{W_{\mathbb{C}}}$ we need some more definitions. We introduce analogs of the counting sequences $(a_{n})$, $\dots$, $(d_{n})$ and generating functions $A(x)$, $\dots$, $D(x)$. Let $h$ be a light weight and $n\in\mathbb{N}_{0}$. We define three sets of $1$-walks of length $n$ in $K_{\mathbb{N}}$: $W_{\mathrm{rec}}(n)$ are the recurrent walks which revisit the start $1$, $W_{b}(n)$ are the walks that end at $1$, and $W_{c}(n)$ are the walks that end at $1$ and the inner avoids $1$. $W(n)$ is of course the set of all $1$-walks of length $n$ in $K_{\mathbb{N}}$. We define $a_{n}^{h}=h(W_{\mathrm{rec}}(n))$, $b_{n}^{h}=h(W_{b}(n))$, $c_{n}^{h}=h(W_{c}(n))$ and $d_{n}^{h}=h(W(n))$. These numbers are correctly defined because $h$ is light and all four series are subseries of $\sum_{w\in W(n)}h(w)$. We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$. We define the generating functions
${\textstyle A_{h}(x)=\sum_{n=0}^{\infty}a_{n}^{h}x^{n},\ B_{h}(x)=\sum_{n=0}^{\infty}b_{n}^{h}x^{n},\ C_{h}(x)=\sum_{n=0}^{\infty}c_{n}^{h}x^{n}\,\text{ and }}$
${\textstyle D_{h}(x)=\sum_{n=0}^{\infty}d_{n}^{h}x^{n}\,.}$
All four are FPS in $\mathbb{C}[[x]]$. The main quantities we want to determine are $a_{n}^{h}$ and $A_{h}(x)$. We generalize Proposition 3.1.
Proposition 5.9
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and the generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be as above. The following formal relations hold in $\mathbb{C}[[x]]$.
  1. 1.
    $A_{h}(x)=C_{h}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=\frac{1}{1-C_{h}(x)}=\sum_{j=0}^{\infty}C_{h}^{j}(x)$, equivalently $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
Proof. 1. Let $n\in\mathbb{N}_{0}$. We need to show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$. For $n\leq 1$ it holds trivially because then $a_{n}^{h}=c_{n}^{h}=0$. For $n\geq 2$ we have $W_{\mathrm{rec}}(n)\neq\emptyset$, this set is in fact countable, and we use the bijection
${\textstyle F\colon W_{\mathrm{rec}}(n)\to\bigcup_{j=0}^{n}W_{c}(j)\times W(n-j)}$
that is defined as follows. For $w=\langle u_{0},\dots,u_{j},\dots,u_{n}\rangle$ in $W_{\mathrm{rec}}(n)$, where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for $i\in[j-1]$, we set $F(w)=(w_{1},w_{2})$, where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. The map $F$ is weight-preserving, or WP, in the sense that for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=(w_{1},w_{2})$ we have $h(w)=h(w_{1})h(w_{2})$. We formally extend $h$ to pairs of walks by setting $h((w,w^{\prime}))=h(w)h(w^{\prime})$. Let $h$ be a light weight. We have, with the help of the bijection $F$ and the above propositions, that
$a_{n}^{h}$ $=$ $h(W_{\mathrm{rec}}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{\mathrm{rec}}(n)])$
[…]
2. Let $n\in\mathbb{N}_{0}$. We need to show the equality
${\textstyle b_{n}^{h}=\sum_{j=0}^{\infty}\sum_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}c_{n_{1}}^{h}c_{n_{2}}^{h}\dots c_{n_{j}}^{h}\,,}$
where the empty sum is set to 00, and the empty product to $1$. For every $n$ the double sum on the right-hand side has only finitely many summands. This equality again holds trivially for $n\leq 1$. If $n=0$ then $b_{0}^{h}=1$ and this $1$ is provided by the summand on the right-hand side with $j=0$ (for $n>0$ this summand is 00). If $n=1$ then $b_{1}^{h}=c_{1}^{h}=0$. For $n\geq 2$ we have $W_{b}(n)\neq\emptyset$, this set is again countable, and we use the bijection
${\textstyle F\colon W_{b}(n)\to\bigcup_{j=1}^{\infty}\bigcup_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}W_{c}(n_{1})\times\cdots\times W_{c}(n_{j})}$
that is defined as follows. For
$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in W_{b}(n)\,,$
where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set $F(w)=\langle w_{1},w_{2},\dots,w_{j}\rangle$ where
$w_{i}=\langle u_{m_{i-1}},\,u_{m_{i-1}+1},\,\dots,\,u_{m_{i}}\rangle,\ i\in[j]$
(thus $n_{i}=m_{i}-m_{i-1}$). Again $F$ is weight-preserving, or WP. We formally extend any weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ to tuples of walks by
${\textstyle h(\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle)=\prod_{i=1}^{j}h(w_{i})\,.}$
Let $h$ be a light weight. We have, with the help of the bijection $F$ and the above propositions, that
$b_{n}^{h}$ $=$ $h(W_{b}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{b}(n)])$
[…]
$\Box$
This proof utilizes our semi-formal approach, or SFA, to generating functions. SFA extends the symbolic method in (enumerative) combinatorics, or SMC, [3, 8, 10, 12, 28] from finite to countable sets. For example, the product formula for generating functions in SMC is based on the identities
$|A\times B|=|A|\cdot|B|\,\text{ and }\,|C\cup D|=|C|+|D|$
for cardinalities of finite sets $A$, $B$, $C$ and $D$, where $C\cap D=\emptyset$. For a formal statement see, for example, [12, 2.2.14]. SFA extends it to identities like the above $A_{h}(x)=C_{h}(x)D_{h}(x)$. The product formula for generating functions in SFA is based on Propositions 5.6 and 5.5 which deal with at most countable sets. We plan to explain SFA in [15].
We say that a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is convex if for every vertex $u\in V(h)$ we have the sum
${\textstyle\sum_{v\in V(h)\setminus\{u\}}h(\{u,\,v\})=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=1\,.}$
We show that these sums exist. If they are all $\geq 1$, we call $h$ superconvex.
Proposition 5.10
Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight and that $u\in V(h)$. Then the series $\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})$ absolutely converges.
Proof. Let $h$ and $u$ be as stated. There is a $1$-walk $w_{0}=\langle u_{0},u_{1},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{n}=u$. We set
${\textstyle c=h(w_{0})=\prod_{i=1}^{n}h(\{u_{i-1},u_{i}\})\ \ (\neq 0)}$
and denote by $W$ the set of $1$-walks in $K_{\mathbb{N}}$ of the form
$\langle u_{0},\,u_{1},\,\dots,\,u_{n},\,u_{n+1}\rangle\,.$
Thus $u_{0}$, $u_{1}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. Then $\sum_{w\in W}h(w)$ absolutely converges because it is a subseries of the absolutely convergent series $\sum_{w\in W(n+1)}h(w)$. By Proposition 5.3 the series
${\textstyle\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})=\frac{1}{c}\sum_{w\in W}h(w)}$
absolutely converges. $\Box$
Convexity is a remnant of the probabilistic approach. It enables us to generalize the identity $D(x)=\frac{1}{1-x}$.
Proposition 5.11
For every convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
${\textstyle D_{h}(x)=\sum_{n=0}^{\infty}h(W(n))x^{n}=\sum_{n=0}^{\infty}x^{n}=\frac{1}{1-x}}\,.$
Proof. Let $h$ be a convex light weight. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=d_{n}^{h}=1$. For $n=0$ this holds as $W(0)=\{(1)\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk. Using Propositions 5.5 and 5.3, the convexity and induction we get that indeed
$h(W(n))$ $=$ ${\textstyle\sum_{w\in W(n)}h(w)}$
$=$ ${\textstyle\sum_{\begin{subarray}{c}w^{\prime}\in W(n-1)\\ \ell(w^{\prime})\in V(h)\end{subarray}}h(w^{\prime})\sum_{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}}h(\{\ell(w^{\prime}),\,v\})}$
$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot 1=1\,.}$
We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because the walks $w^{\prime}$ not satisfying it have zero weights. $\Box$
We generalize Definition 3.2.
Definition 5.12 ($U(R)$ 2)
Let $R>0$ be a real number and $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$ be a FPS in $\mathbb{C}[[x]]$. We define
${\textstyle U(R)=\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}R^{j}=\sum_{n=0}^{\infty}u_{j}R^{j}}\ \ (\in\mathbb{C})\,,$
if this limit exists..
This definition is slightly more restrictive than Definition 3.2 because we do not allow the value $U(R)=+\infty$.
Proposition 5.13
Suppose that $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$ and $V(x)=\sum_{n=0}^{\infty}v_{n}x^{n}$ are FPS in $\mathbb{C}[[x]]$, that the sum $V(1)$ exists and that formally $U(x)=V(x)\frac{1}{1-x}$. Then
$\lim_{n\to\infty}u_{n}=\lim_{n\to\infty}(v_{0}+v_{1}+\dots+v_{n})=V(1)\ \ (\in\mathbb{C})\,.$
Proof. Again, $u_{n}=v_{0}+v_{1}+\dots+v_{n}$ for every $n\in\mathbb{N}_{0}$. $\Box$
We make use of the next classical theorem, see [29, Chapter II.7]. For a real number $R>0$ we denote by
$K_{R}=\{x\in\mathbb{C}:\;|x|<R\}$
the open disc in $\mathbb{C}$ with radius $R$ and center at 00. It is well known that if $U(x)$ is in $\mathbb{C}[[x]]$ and $U(R)$ exists, then the series $U(x)$ absolutely converges for every $x\in K_{R}$.
Theorem 5.14 (Abel’s)
Let $R>0$ and $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$ be a FPS in $\mathbb{C}[[x]]$ such that the sum $U(R)$ exists. Then the sum $F_{U}\colon K_{R}\to\mathbb{C}$ of $U(x)$ has the limit
$\lim_{\begin{subarray}{c}x\to R\\ x\in[0,\,R)\end{subarray}}F_{U}(x)=U(R)\ \ (\in\mathbb{C})\,.$
Corollary 5.15
Suppose that $U(x)=\sum_{n=0}^{\infty}u_{n}x^{n}$, $V(x)=\sum_{n=0}^{\infty}v_{n}x^{n}$ and $W(x)=\sum_{n=0}^{\infty}w_{n}x^{n}$ are FPS in $\mathbb{C}[[x]]$ such that formally $U(x)=V(x)W(x)$. If the sums $U(1)$, $V(1)$ and $W(1)$ exist, then
$U(1)=V(1)W(1)\ \ (\in\mathbb{C})\,.$
Proof. Propositions 5.5 and 5.6 imply that $F_{U}(x)=F_{V}(x)F_{W}(x)$ for every $x\in K_{1}$. Thus by Theorem 5.14,
[…]
$\Box$
In contrast to the nonnegative real case mentioned in the remark after Corollary 3.6, in the complex case one cannot get the equality $U(1)=V(1)W(1)$ without Abel’s theorem only by manipulating series. The reason is that the involved series need not be absolutely convergent.
Corollary 5.16
Suppose that $U(x)$ and $V(x)$ are FPS in $\mathbb{C}[[x]]$ such that formally $V(x)=1-\frac{1}{U(x)}$. If the sums $U(1)$ and $V(1)$ exist, then $U(1)\neq 0$ and
${\textstyle V(1)=1-\frac{1}{U(1)}\ \ (\in\mathbb{C})\,.}$
Proof. We have formally $V(x)U(x)=U(x)-1$ and argue, using Theorem 5.14, as in the previous corollary. $\Box$
Corollary 5.17
Suppose that $U(x)$, $V(x)$ and $W(x)$ are FPS in $\mathbb{C}[[x]]$ such that formally $W(x)=V(x)+W(x)\cdot U(x)^{2}$. If the sums $U(1)$, $V(1)$ and $W(1)$ exist, then
$W(1)=V(1)+W(1)\cdot U(1)^{2}\ \ (\in\mathbb{C})\,.$
Proof. We argue as in the previous proof. $\Box$
[…]
Theorem 5.18 (general $1$-recurrence)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Using the above notation we get the following.
  1. 1.
    (In this case $D_{h}(1)$ exists.) Suppose that the sums $D_{h}(1)$, $C_{h}(1)$, $B_{h}(1)$ and $A_{h}(1)$ exist — a sufficient condition for it is $D_{|h|}(1)<+\infty$. Then $B_{h}(1)\neq 0$ and
    $A_{h}(1)={\textstyle\sum_{n=0}^{\infty}\sum_{w\in W_{\mathrm{rec}}(n)}h(w)=\big{(}1-\frac{1}{B_{h}(1)}\big{)}D_{h}(1)\ \ (\in\mathbb{C})\,.}$
  2. 2.
    (In this case $D_{h}(1)$ does not exist.) Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ and $D_{h}(1)=+\infty$. Then $c_{2}^{h}>0$ and $a_{n+2}^{h}\geq c_{2}^{h}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence $A_{h}(1)=+\infty$.
Proof. 1. If $D_{|h|}(1)<+\infty$ then the four series with the sums $D_{h}(1)$, $C_{h}(1)$, $B_{h}(1)$ and $A_{h}(1)$ absolutely converge. This follows from the fact that for every $n\in\mathbb{N}_{0}$,
$|d_{n}^{h}|,\,|c_{n}^{h}|,\,|b_{n}^{h}|,\,|a_{n}^{h}|\leq d_{n}^{|h|}\,.$
These inequalities follow from Proposition 5.1 and the fact that for every $n$ in $\mathbb{N}_{0}$ the complex series $\sum_{w\in W(n)}h(w)$, $\sum_{w\in W_{c}(n)}h(w)$, $\sum_{w\in W_{b}(n)}h(w)$ and $\sum_{w\in W_{a}(n)}h(w)$ are subseries of $\sum_{w\in W(n)}h(w)$.
Suppose that $D_{h}(1)$, $C_{h}(1)$, $B_{h}(1)$ and $A_{h}(1)$ exist. By part 1 of Proposition 5.9 we have formally $A_{h}(x)=C_{h}(x)D_{h}(x)$. By Corollary 5.15, $A_{h}(1)=C_{h}(1)D_{h}(1)$ ($\in\mathbb{C}$). By part 2 of Proposition 5.9 and by Corollary 5.16, $B_{h}(1)\neq 0$ and $C_{h}(1)=1-\frac{1}{B_{h}(1)}$. The stated formula follows.
2. Suppose that $h$ is nonnegative and $D_{h}(1)=+\infty$. Thus $d_{n}^{h}>0$ for some $n>0$ and there must be a vertex $u\in\mathbb{N}\setminus\{1\}$ with $h(\{1,u\})>0$. The walk $\langle 1,u,1\rangle$ then shows that $c_{2}^{h}>0$. Since by part 1 of Proposition 5.9 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$, the stated inequality follows. Hence $A_{h}(1)=+\infty$. $\Box$
We remark that we cannot completely characterize either case of the theorem. The same holds for the remaining three main theorems in this section.
The first of them is another extension of Theorem 3.9 on $\overline{0}$-recurrence to $\mathrm{W_{\mathbb{C}}}$ that employs convex weights. We divide the theorem in two cases according to the existence of $B_{h}(1)$.
Theorem 5.19 (convex $1$-recurrence)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be any convex light weight. Using the above notation we get the following.
  1. 1.
    (In this case $B_{h}(1)$ exists.) Suppose that the sums $B_{h}(1)$ and $C_{h}(1)$ exist — a sufficient condition for it is $B_{|h|}(1)<+\infty$. Then $B_{h}(1)\neq 0$ and
    $\lim_{n\to\infty}a_{n}^{h}=\lim_{n\to\infty}h(W_{\mathrm{rec}}(n))={\textstyle 1-\frac{1}{B_{h}(1)}\ \ (\in\mathbb{C})\,.}$
  2. 2.
    (In this case $B_{h}(1)$ does not exist.) Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ and $B_{h}(1)=+\infty$. Then
    $\lim_{n\to\infty}a_{n}^{h}=\lim_{n\to\infty}{\textstyle\sum_{w\in W_{\mathrm{rec}}(n)}h(w)}=1\,.$
Proof. 1. If $B_{|h|}(1)<+\infty$ then we use the bounds $|b_{n}^{h}|,|c_{n}^{h}|\leq b_{n}^{|h|}$ and proceed as in the proof of part 1 of the previous theorem.
Suppose that $B_{h}(1)$ and $C_{h}(1)$ exist. By part 1 of Proposition 5.9, Proposition 5.11 and by Proposition 5.13, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)$. By part 2 of Proposition 5.9 and Corollary 5.16, $B_{h}(1)\neq 0$ and $C_{h}(1)=1-\frac{1}{B_{h}(1)}$. The stated formula follows.
2. Suppose that $h$ is nonnegative and $B_{h}(1)=+\infty$. By Proposition 5.11, $D_{h}(x)=\frac{1}{1-x}$. Thus all coefficients $a_{n}^{h}$, $b_{n}^{h}$ and $c_{n}^{h}$ lie in $[0,1]$ and formal power series $A_{h}(x)$, $B_{h}(x)$ and $C_{h}(x)$ converge on $[0,1)$. By part 1 of Proposition 5.9, formally $A_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 3.3, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)$. By part 2 of Proposition 5.9 and Corollary 3.5, $C_{h}(1)=1$. The value $1$ of the limit follows. $\Box$
Unlike below in Theorems 5.22 and 5.23 on $v$-recurrence, we do not restrict weights by any vertex-transitivity. This is interesting since we generalized the case of vertex-transitive graphs $G_{d}$.
We generalize Pólya’s claim that “The probability obviously grows with $n$.” Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $w=\langle u_{0},u_{1}\dots,u_{n}\rangle$. We define
$W(v,\,n)=\{w\in W(n):\;\exists\,i\in[n]\,(u_{i}=v)\}\;.$
It is the set of $v$-recurrent $1$-walks of length $n$ in $K_{\mathbb{N}}$.
Proposition 5.20
If $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a superconvex light weight, then
$h(W(v,0))\leq h(W(v,1))\leq h(W(v,2))\leq\cdots$
for every $v\in\mathbb{N}$.
Proof. Let $v\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$. Then (recall that $\ell(w)$ is the last vertex of $w$)
$h(W(v,\,n))$ $=$ ${\textstyle\sum_{w\in W(v,\,n)}h(w)=\sum_{\begin{subarray}{c}w\in W(v,\,n)\\ \ell(w)\in V(h)\end{subarray}}h(w)\cdot 1}$
$\leq$ ${\textstyle\sum_{\begin{subarray}{c}w\in W(v,\,n)\\ \ell(w)\in V(h)\end{subarray}}h(w)\cdot\sum_{u\in\mathbb{N}\setminus\{\ell(w)\}}h(\{\ell(w),\,u\})}$
[…]
On the second line we used the superconvexity of $h$. We can add and remove the condition $\ell(w)\in V(h)$ because walks not satisfying it have zero weight. $\Box$
We introduce quantities needed for generalizing $\overline{v}$-recurrence. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight, $v\in\mathbb{N}\setminus\{1\}$ and $n\in\mathbb{N}_{0}$. We define sequences of weights of sets of $1$-walks of length $n$ in $K_{\mathbb{N}}$. Namely, $(a_{n}^{h})^{\prime}=h(W(v,n))$, $b_{n}^{h}=h(W_{b}(n))$ is as before, $(b_{n}^{h})^{\prime}=h(W_{b^{\prime}}(n))$ is the weight of walks ending at $1$ and avoiding $v$, $(c_{n}^{h})^{\prime}=h(W_{c^{\prime}}(n))$ is the weight of walks ending at $v$ and with the inner avoiding $v$, $(c_{n}^{h})^{\prime\prime}=h(W_{c^{\prime\prime}}(n))$ is the weight of walks ending at $v$ and with the inner avoiding both $1$ and $v$, and $d_{n}^{h}=h(W(n))$ is as before. We have $(a_{0}^{h})^{\prime}=(c_{0}^{h})^{\prime}=(c_{0}^{h})^{\prime\prime}=0$ and $b_{0}^{h}=(b_{0}^{h})^{\prime}=d_{0}^{h}=1$. Besides the previous generating functions $B_{h}(x)$ and $D_{h}(x)$ we consider
${\textstyle A_{0,h}(x)=\sum_{n=0}^{\infty}(a_{n}^{h})^{\prime}x^{n},\ B_{0,h}(x)=\sum_{n=0}^{\infty}(b_{n}^{h})^{\prime}x^{n},}$
${\textstyle C_{0,h}(x)=\sum_{n=0}^{\infty}(c_{n}^{h})^{\prime}x^{n}\,\text{ and }\,C_{1,h}(x)=\sum_{n=0}^{\infty}(c_{n}^{h})^{\prime\prime}x^{n}\,.}$
All are in $\mathbb{C}[[x]]$. The dependence on $v$ is left implicit in the notation.
Let $v\in\mathbb{N}$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and for every edge $\{a,b\}\in\mathbb{N}_{2}$ we have $h(\{a,b\})=h(\{f(a),f(b)\})$. We generalize Proposition 3.10.
Proposition 5.21
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and let the generating functions $A_{0,h}(x)$, $B_{h}(x)$, $B_{0,h}(x)$, $C_{0,h}(x)$, $C_{1,h}(x)$ and $D_{h}(x)$ be as above. The following formal relations hold.
  1. 1.
    $A_{0,h}(x)=C_{0,h}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=B_{0,h}(x)+C_{0,h}(x)^{2}B_{h}(x)$, equivalently $B_{h}(x)=\frac{B_{0,h}(x)}{1-C_{0,h}(x)^{2}}$.
  3. 3.
    $C_{0,h}(x)=B_{0,h}(x)C_{1,h}(x)$.
Proof. We give less details compared to the proof of Proposition 5.9. We argue on the high level of SFA, the semi-formal approach to generating functions.
1. Any walk $w$ in $\bigcup_{n=0}^{\infty}W(v,n)$ uniquely splits at the first visit of $v$ as $w=w_{1}w_{2}$, where the walks $w_{1}$ are weight-counted by $C_{0,h}(x)$ and $w_{2}$ are arbitrary $v$-walks. By the $v$-transitivity of $h$, the walks $w_{2}$ are weight-counted by $D_{h}(x)$. The first relation follows.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts $1$-walks $w$ in $K_{\mathbb{N}}$ ending at $1$. Those avoiding $v$ are weight-counted by $B_{0,h}(x)$. Those visiting $v$ split uniquely as $w=w_{1}w_{2}w_{3}$, where the walk $w_{1}$ starts at $1$, ends at $v$ and the inner avoids $v$, the walk $w_{2}$ starts and ends at $v$, and the walk $w_{3}$ starts at $v$, ends at $1$ and the inner avoids $v$. By reversing the walk $w_{3}$ we see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{0,h}(x)$. By the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$. The second relation follows.
3. We consider walks $w$ weight-counted by $C_{0,h}(x)$. They start at $1$, end at $v$ and the inner avoids $v$. They uniquely split at the last visit of $1$ in two walks as $w=w_{1}w_{2}$, where $w_{1}$ and $w_{2}$ are weight-counted by $B_{0,h}(x)$ and $C_{1,h}(x)$, respectively. The third relation follows. $\Box$
A main result is the first extension of Theorem 3.11 on $\overline{v}$-recurrence to $\mathrm{W_{\mathbb{C}}}$. It concerns general weights $h:\mathbb{N}_{2}\to\mathbb{C}$. Recall that for any weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we denote by $V(h)$ ($\subset\mathbb{N}$) the vertex set of the component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. For $x\in\mathbb{C}$ we define
$\mathrm{sqrt}(x)=\{y\in\mathbb{C}\colon\;y^{2}=x\}\,.$
Always $|\mathrm{sqrt}(x)|=2$, except for $\mathrm{sqrt}(0)=\{0\}$. For two numbers $x,z\in\mathbb{C}$ we define
$\mathrm{sqrt}(x)\cdot z=\{zy\colon\;y\in\mathrm{sqrt}(x)\}\,.$
As before we divide the theorem in two cases according to the existence of $D_{h}(1)$, but now we also have the singular case when $v\not\in V(h)$.
Theorem 5.22 (general $v$-recurrence)
Suppose that $v\in\mathbb{N}\setminus\{1\}$ and that $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight. Using the above notation we get the following.
  1. 1.
    If $v\not\in V(h)$ then $(a_{n}^{h})^{\prime}=h(W(v,n))=0$ for every $n\in\mathbb{N}_{0}$.
  2. 2.
    (In this case $D_{h}(1)$ exists.) Suppose that the sums $D_{h}(1)$, $C_{0,h}(1)$, $B_{0,h}(1)$, $B_{h}(1)$ and $A_{0,h}(1)$ exist — a sufficient condition for it is $D_{|h|}(1)<+\infty$. Then $B_{h}(1)\neq 0$ and
    $A_{0,h}(1)={\textstyle\sum_{n=0}^{\infty}\sum_{w\in W(v,\,n)}h(w)}\in{\textstyle\mathrm{sqrt}\big{(}1-\frac{B_{0,\,h}(1)}{B_{h}(1)}}\big{)}\cdot D_{h}(1)\ \ (\subset\mathbb{C})\,.$
  3. 3.
    (In this case $D_{h}(1)$ does not exist and $v\in V(h)$.) Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$, $D_{h}(1)=+\infty$ and $v\in V(h)$. Then for an $m\in\mathbb{N}$ we have $(c_{m}^{h})^{\prime}>0$ and $(a_{n+m}^{h})^{\prime}\geq(c_{m}^{h})^{\prime}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence $A_{0,h}(1)=+\infty$.
Proof. Let $v$ and $h$ be as stated.
1. Every walk in $K_{\mathbb{N}}$ joining $1$ and $v$ has weight 00, which implies that $(a_{n}^{h})^{\prime}=0$ for every $n\in\mathbb{N}_{0}$.
2. For $D_{|h|}(1)<+\infty$ we proceed as in the proof of part 1 of Theorem 5.18. We have bounds
$|d_{n}^{h}|,\,|(c_{n}^{h})^{\prime}|,\,|(b_{n}^{h})^{\prime}|,\,|b_{n}^{h}|,\,|(a_{n}^{h})^{\prime}|\leq d_{n}^{|h|}\,.$
Suppose that $D_{h}(1)$, $C_{0,h}(1)$, $B_{0,h}(1)$, $B_{h}(1)$ and $A_{0,h}(1)$ exist. By part 2 of Proposition 5.21 and by Corollary 5.17, $B_{h}(1)=B_{0,h}(1)+C_{0,h}(1)^{2}B_{h}(1)$. From the proof of part 1 of Theorem 5.18 we know that $B_{h}(1)\neq 0$. Hence $C_{0,h}(1)\in\mathrm{sqrt}\big{(}1-\frac{B_{0,h}(1)}{B_{h}(1)}\big{)}$. By part 1 of Proposition 5.21, formally $A_{0,h}(x)=C_{0,h}(x)D_{h}(x)$. By Corollary 5.15, $A_{0,h}(1)=C_{0,h}(1)D_{h}(1)$ and the stated formula follows.
3. Suppose that $h$ is nonnegative, $D_{h}(1)=+\infty$ and $v\in V(h)$. The shortest walk of length $m$ joining $1$ and $v$ in $G(h)$ shows that $(c_{m}^{h})^{\prime}>0$. Since by part 1 of Proposition 5.21 we have $(a_{n}^{h})^{\prime}=\sum_{j=0}^{n}(c_{j}^{h})^{\prime}d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$, the stated inequality follows. Hence $A_{0,h}(1)=+\infty$. $\Box$
The last main result is the second extension of Theorem 3.11 on $\overline{v}$-recurrence to $\mathrm{W_{\mathbb{C}}}$. It employs convex weights. We again divide the theorem in two parts according to the existence of $B_{h}(1)$, and have the singular case when $v\not\in V(h)$.
Theorem 5.23 (convex $v$-recurrence)
Suppose that $v\in\mathbb{N}\setminus\{1\}$ and that $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a convex $v$-transitive light weight. Using the above notation we get the following.
  1. 1.
    If $v\not\in V(h)$ then $(a_{n}^{h})^{\prime}=h(W(v,n))=0$ for every $n\in\mathbb{N}_{0}$.
  2. 2.
    (In this case $B_{h}(1)$ exists.) Suppose that the sums $B_{h}(1)$, $B_{0,h}(1)$ and $C_{0,h}(1)$ exist — a sufficient condition for it is $D_{|h|}(1)<+\infty$. Then $B_{h}(1)\neq 0$ and
    $\lim_{n\to\infty}(a_{n}^{h})^{\prime}=\lim_{n\to\infty}h(W(v,\,n))\in{\textstyle\mathrm{sqrt}\big{(}1-\frac{B_{0,\,h}(1)}{B_{h}(1)}}\big{)}\ \ (\subset\mathbb{C})\,.$
  3. 3.
    (In this case $B_{h}(1)$ does not exist and $v\in V(h)$.) Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$, $B_{h}(1)=+\infty$ and $v\in V(h)$. Then
    $\lim_{n\to\infty}(a_{n}^{h})^{\prime}=\lim_{n\to\infty}{\textstyle\sum_{w\in W(v,\,n)}h(w)}=1\,.$
Proof. Let $v$ and $h$ be as stated.
1. Again $(a_{n}^{h})^{\prime}=0$ for every $n\in\mathbb{N}_{0}$ because $v$ is unreachable from $1$ in $G(h)$.
2. If $D_{|h|}(1)<+\infty$ then we proceed as in the proof of part 1 of Theorem 5.18; now we have the bounds
$|b_{n}^{h}|,\,|(b_{n}^{h})^{\prime}|,\,|(c_{n}^{h})^{\prime}|\leq d_{n}^{|h|}\,.$
Suppose that $B_{h}(1)$, $B_{0,h}(1)$ and $C_{0,h}(1)$ exist. From part 2 of Proposition 5.21 and Corollary 5.17 we get that
$B_{h}(1)=B_{0,h}(1)+C_{0,h}(1)^{2}B_{h}(1)\,.$
From the proof of part 2 of Theorem 5.18 we know that $B_{h}(1)\neq 0$. Hence $C_{0,h}(1)\in\mathrm{sqrt}\big{(}1-\frac{B_{0,h}(1)}{B_{h}(1)}\big{)}$. By part 1 of Proposition 5.21, Proposition 5.11 and Proposition 5.13, $\lim_{n\to\infty}(a_{n}^{h})^{\prime}=C_{0,h}(1)$. The stated formula follows.
3. We assume that $h$ is nonnegative, $B_{h}(1)=+\infty$ and $v\in V(h)$. By Proposition 5.11, $D_{h}(x)=\frac{1}{1-x}$. Hence for every $n\in\mathbb{N}_{0}$ the coefficients $(a_{n}^{h})^{\prime}$, $b_{n}^{h}$, $(b_{n}^{h})^{\prime}$, $(c_{n}^{h})^{\prime}$ and $(c_{n}^{h})^{\prime\prime}$ lie in $[0,1]$ and we can use the functions
[…]
Since $v\in V(h)$, we have $(c_{n}^{h})^{\prime\prime}>0$ for some $n>0$ (consider the shortest walk in $G(h)$ joining $1$ and $v$). Thus we deduce by the same argument as in the proof of the case $d\leq 2$ of Theorem 3.11 that $F_{B_{0,h}}$ is bounded on $[0,1)$. This boundedness, the assumption that $B_{h}(1)=+\infty$, part 2 of Proposition 5.21 and Theorem 3.4 give that $C_{0,h}(1)=1$. By part 1 of Proposition 5.21, $D_{h}(x)=\frac{1}{1-x}$, and Proposition 3.3 we get the stated limit that $\lim_{n\to\infty}(a_{n}^{h})^{\prime}=C_{0,h}(1)=1$. $\Box$
We extend Corollary 3.12 to $\mathrm{W_{\mathbb{C}}}$.
Corollary 5.24
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive light nonnegative weight such that the FPS $D_{h}(x)$ converges on $[0,1)$, the sum $B_{h}(1)=+\infty$ and $v\in V(h)$. Then
$\lim_{x\to 1}F_{B_{h}}(x)=+\infty\,\text{ and }\,\lim_{x\to 1}{\textstyle\frac{F_{B_{0,h}}(x)}{F_{B_{h}}(x)}=0\,.}$
Proof. The first limit follows from $B_{h}(1)=+\infty$ and Theorem 3.4. We know that on $[0,1)$, $F_{B_{h}}\geq 1$ and $\frac{F_{B_{0,h}}}{F_{B_{h}}}=1-F_{C_{0,h}}^{2}$. Let $x\to 1$. Then $F_{C_{0,h}}\to 1$ by the argument in the proof of the case $d\leq 2$ of Theorem 3.11 (note that $(c_{n}^{h})^{\prime\prime}>0$ for some $n>0$ because $v\in V(h)$) and the second limit follows. $\Box$
6 Concluding remarks
Since this paper is already long, we decided to move several concrete examples in the follow-up paper [16]. There we obtain analogues of the four main theorems in the previous section for weights
$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,x_{2},\,\dots,\,x_{k}]]\,.$
The algebras $\mathbb{C}[[x_{1},x_{2},\dots,x_{k}]]$ of FPS are endowed with non-Archimedean norms. Non-Archimedean infinite series are simpler than the complex ones we dealt with above. Weighting objects with polynomials or formal power series is a basic tool in enumerative combinatorics.